### v3

# Signal-SGN: A Spiking Graph Convolutional Network for Skeleton Action Recognition via Learning Temporal-Frequency Dynamics

Abstract
For skeleton-based action recognition, Graph Convolutional Networks (GCNs) are effective models. Still, their reliance on floating-point computations leads to high energy consumption, limiting their applicability in battery-powered devices. While energy-efficient, Spiking Neural Networks (SNNs) struggle to model skeleton dynamics, leading to suboptimal solutions. We propose Signal-SGN (Spiking Graph Convolutional Network), which utilizes the temporal dimension of skeleton sequences as the spike time steps and represents features as multi-dimensional discrete stochastic signals for temporal-frequency domain feature extraction. It combines the 1D Spiking Graph Convolution (1D-SGC) module and the Frequency Spiking Convolution (FSC) module to extract features from the skeleton represented as spiking form. Additionally, the Multi-Scale Wavelet Transform Feature Fusion (MWTF) module is proposed to extract dynamic spiking features and capture frequency-specific characteristics, enhancing classification performance. Experiments across three large-scale datasets reveal Signal-SGN exceeding state-of-the-art SNN-based methods in accuracy and computational efficiency while attaining comparable performance with GCN methods and significantly reducing theoretical energy consumption.
Index Terms:
Skeleton-based Action Recognition, Spiking Neural Networks, Frequency Domain Transform
I Introduction
Skeleton-based action recognition, as a critical subfield of human action recognition, has made significant strides in recent years. Leveraging sparse representations obtained from pose estimation and sensor-based data, it enables effective classification of human actions [1]. Numerous methods have been proposed for this task, including CNNs [2], LSTMs [3], Transformers [4], and Graph Convolutional Networks (GCNs) [5]. Among them, ST-GCN is a pioneering method that laid the foundation for GCN-based approaches [6]. Several methods,such as 2s-AGCN [7] and Dynamic GCN [8], have been proposed subsequently, further improving accuracy in action recognition. Lightweight models like Shift-GCN [5] and CTR-GCN [9] have further enhanced accuracy while reducing computational costs. However, these methods are all based on artificial neural networks (ANNs), which rely on floating-point operations (FLOPs), leading to significant energy consumption that limits their practicality in energy-constrained applications.
Spiking Neural Networks (SNNs), regarded as third-generation neural networks, utilize spike-form features encoded as 0s and 1s to mimic neural dynamics. Unlike ANNs, SNNs process information in an event-driven manner, enabling significantly lower energy consumption while maintaining efficient computation [10]. Various SNN models have been proposed, leveraging their energy efficiency, and have been applied to a wide range of tasks, including image classification and point cloud estimation [11, 12, 13, 14, 15]. Among these, Spiking Graph Networks (SGNs) combine graph convolution with SNNs, achieving remarkable energy savings. For instance, SGNs reduce energy consumption by nearly 99% compared to vanilla GCNs [16]. Specifically, SGN has been utilized for skeleton action recognition, cutting energy consumption to less than 5% of GCNs while maintaining comparable accuracy [17]. Despite its advantages, [17] treats dynamic skeleton data as static image data, artificially extending it along the spike time step dimension. This approach neglects the time-sensitive computational capabilities of SNNs, failing to effectively capture the inherent temporal dynamics of skeleton data. As a result, it increases computational complexity, hampers performance, and poses significant challenges for efficient training on standard deep-learning hardware. Furthermore, SGNs, like other SNN-based models [13, 14, 15], perform a simplistic Global Average Pooling (GAP) over the spike time step dimension in the task head, failing to thoroughly analyze the discrete and aperiodic nature of spike-form features. This oversimplification limits the model’s ability to capture temporal patterns, reducing its task performance.
To overcome these challenges, we propose Signal-SGN for skeleton action recognition, utilizing the temporal dimension of skeleton sequences as the spiking time step and representing skeleton data as spiking-form features with multi-dimensional discrete stochastic properties. The architecture comprises two primary modules: a one-dimensional Spiking Graph Convolution (1D-SGC) module and a Frequency Spiking Convolution (FSC) module. The 1D-SGC extracts feature from single-frame skeleton graphs within temporal sequences and represents them as the spiking-form. Meanwhile, the FSC utilizes the Fast Fourier Transform (FFT) to transform the spiking-form features of each frame into frequency-domain and leverages complex spiking convolution for advanced feature extraction. Through successive 1D-SGC-FSC layers, the features remain discrete and aperiodic. We formulate a Multi-Scale Wavelet Transform Feature Fusion (MWTF) module to address the limitations of simple GAP, which fails to effectively capture the temporal dynamics and frequency-specific characteristics of spiking-form features. The module exploits discrete wavelet transform (DWT) to decompose features into distinct frequency components and integrates a spiking cross-attention mechanism for robust feature fusion. The proposed module significantly improves classification performance. Our main contributions can be summarized as follows:
  • We propose a novel Signal-SGN network comprising multi-layer 1D-SGC-FSC that captures temporal-frequency domain spatial features of single-frame skeleton graphs and leverages the time-sensitive computational properties of SNNs to effectively model sequence-level temporal dynamics.
  • We propose the MWTF module, which achieves multi-scale frequency-domain decomposition and interaction between frequency-specific features to extract dynamic spiking features, capture frequency-specific characteristics, and improve classification performance.
  • Extensive experiments across three large-scale demonstrate that Signal-SGN surpasses State-of-the-Art (SOTA) SNN methods in terms of computational costs, and accuracy while achieving results comparable to GCN-based methods with significantly lower energy consumption.
II Related Work
II-A Skeleton-Based Action Recognition.
Deep learning methods have played a significant role in skeleton-based action recognition. Early approaches in this field achieved significant success using CNNs[2] and RNNs[3]. Subsequent research has primarily focused on using Transformers[4] and GCNs[6, 7, 5, 18, 9]. These numerous GCN-based methods typically involve nearly multi-layer stacks of Graph Convolutional Networks and Temporal Convolutional Networks. However, GCN-based networks are intrinsically tied to ANN architectures, which rely on floating-point computations. Such reliance leads to substantial energy consumption and fundamentally misaligns with practical requirements for energy-efficient solutions.
II-B Spiking Neural Networks.
SNNs, inspired by biological neural networks, offer an energy-efficient alternative to traditional neural networks [10]. They can be implemented through two primary approaches: ANN-to-SNN conversion [19] and direct training using algorithms like STDP [20] or surrogate gradients [21]. The fundamental units, such as Leaky Integrate-and-Fire (LIF) neurons, are notable for their simplicity and biological plausibility [22]. The differential equation describes the dynamics of the LIF neuron:
$\tau_{m}\frac{dV(t)}{dt}=-V(t)+R\cdot I(t),$ (1)
where $V(t)$ is the membrane potential, $\tau_{m}$ is the time constant indicating potential decay, $R$ is the membrane resistance, and $I(t)$ is the input current. When $V(t)$ reaches the threshold $V_{\text{th}}$, the neuron fires a spike and resets to $V_{\text{reset}}$. The LIF neuron model accurately replicates the spiking dynamics characteristic of neurons within SNNs.
Inspired by ANN architectures, corresponding SNN models have been proposed and applied to tasks such as image classification, event-driven processing, and EEG analysis[23, 24, 25]. Among these, various spiking-based Transformers have played significant roles[13, 14, 15]. Meanwhile, [17] introduces Spiking Graph Networks (SGNs) for skeleton-based action recognition for the first time, achieving promising results. However, [17] treats the skeleton sequence data as static and artificially extends them along the spike time step dimension, which increases storage and computational demands. Additionally, SGNs fail to capture the discrete and non-periodic nature of spike-form features and rely on simple GAP to reduce the spike time step dimension before the classification head, limiting classification performance and accuracy.
II-C Frequency Domain Analysis.
In neural network design, frequency domain analysis is crucial for improving performance and robustness. The FFT and Wavelet Transform (WT) enhance feature extraction within CNNs for applications such as image compression [26]. FFT also plays an important role in sequential data analysis within RNNs, which is particularly beneficial for applications such as speech recognition [27]. GCNs integrate Fourier transforms (FT) to analyze graph signals [28]. Meanwhile, SNNs combined with WT enable effective spatial-frequency domain learning, achieving energy efficiency and high performance [29]. In the context of SNNs, skeleton sequences are represented as spiking-form feature, resembling discrete multi-dimensional stochastic signals. Exploring their frequency domain characteristics has significant potential for the analysis of skeleton action recognition within SNNs. However, research in the area is currently insufficient.
Figure 1: The overview of the proposed Signal-SGN method.
III Method
This section presents the overall architecture of Signal-SGN, provided in Figure 1. It begins with a description of the backbone network, including the 1D-SGC module and the FSC module, which constitute the core components of the model. Following this, the MWTF module is explained, emphasizing its role for enhancing performance.
III-A The Backbone of Signal-SGN
As illustrated in Figure 1, Signal-SGN processes three-dimensional skeleton data $X\in\mathbb{R}^{C\times T\times V}$, where $C$ represents spatial coordinates $(x,y,z)$, $T$ denotes the number of frames and spike time steps, and $V$ is the number of joints per frame. The spatial coordinates, which contain both positive and negative values, pose challenges to SNNs due to rate coding and binarization, leading to information loss. To address this, the coordinates are normalized to $[0,1]$ using the formula $a^{\prime}=(a-a_{\min})/(a_{\max}-a_{\min})$, where $a\in\{x,y,z\}$. After normalization, the skeleton data is transformed into $X_{\text{i}}\in\mathbb{R}^{T\times C^{\prime}\times V}$, where $C^{\prime}$ corresponds to the normalized three-dimensional coordinates.
In the 1D-SGC module, inspired by GCN, we integrate topological information by normalizing adjacency-based weights to aggregate attributes from neighboring nodes, as shown in Figure 1(a). The skeleton can be represented as a graph $\mathcal{G}(V,E)$, where the joints form a set of vertices $V$ and the bones represent the edges $E$. This structure can be expressed as an adjacency matrix $A\in\mathbb{R}^{V\times V}$, where $A_{i,j}=1$ if joints $i$ and $j$ are physically connected, and $A_{i,j}=0$ otherwise. To incorporate topological relationships into computations, the normalized adjacency matrix $S$ is computed as follows:
$S=\Delta^{-\frac{1}{2}}\widetilde{A}\Delta^{-\frac{1}{2}},$ (2)
where $\widetilde{A}=A+I$ is the adjacency matrix with added self-connections, and $\Delta$ is the degree matrix of $\widetilde{A}$. This normalization ensures that the influence of each node is appropriately scaled based on its connectivity, facilitating the effective integration of topological information into the model. Subsequently, we apply one-dimensional convolution (Conv1d) and one-dimensional Batch Normalization(BN) along the $C^{\prime}$ dimension to extract features for each skeleton node within each spike time step, effectively aggregating features across the skeleton joints dimension.
$G_{\text{i}}=\mathrm{BN}\left[\sum_{n=0}^{N}\operatorname{Conv1d}\left(X_{\text{i}}*S^{n}\right)\right],$ (3)
in which $S^{n}$ represents a set of matrices decomposed from $S$, $N$ refers to the number of different relational matrices in the decomposition, and * denotes matrix multiplication. The resulting output $G_{\text{i}}\in\mathbb{R}^{T\times D\times V}$, where $D$ represents the hidden channel dimension.
[…]
$G^{(0)}_{\text{o}}=SN\left(G_{\text{i}}\right)+\Phi\left(X_{\text{i}}\right).$ (4)
Building upon the approach proposed in [13], we integrate the spike self-attention (SSA) mechanism after the graph convolution operation to enhance feature representation. The formula is as follows:
$Q^{(0)}=\Phi_{Q}(G^{(0)}_{\text{o}}),K^{(0)}=\Phi_{K}(G^{(0)}_{\text{o}}),V^{(0)}=\Phi_{V}(G^{(0)}_{\text{o}}),$ (5)
$G^{(0)}=G^{(0)}_{\text{o}}+\Phi\left(Q^{(0)}K^{(0)T}V^{(0)}*s\right),$ (6)
where $s$ is the scaling factor to control the large value of the matrix multiplication result.
Following the 1D-SGC module, the features are re-transformed by $SN$ into spiking-form feature—discrete, aperiodic signal sequences encoded as binary values (0s and 1s). To further extract meaningful frequency-domain feature, we propose the FSC module, as shown in Figure 1(b). This module first leverages a learnable window function applied along the joint dimension $V$ to mitigate spectral leakage, enhancing the precision of frequency-domain analysis and improving feature extraction robustness, defined as:
$w_{V}(v)=\alpha-(1-\alpha)\cdot\cos\left(\frac{2\pi v}{V-1}\right),\quad 0\leq v\leq V-1.$ (7)
where $v$ is the joint index and $\alpha$ is a learnable parameter controlling the window shape. The learnability of $\alpha$ allows the module to adapt the window function to specific characteristics. The windowed features are then processed as:
$F^{(0)}_{\text{i}}=w_{V}(v)\cdot SN(G^{(0)})\in\mathbb{R}^{T\times D\times V}.$ (8)
We apply FFT along the $V$ dimension to transform the features from the temporal domain to the frequency domain:
$F^{(0)}_{\text{o}}=\sum_{v=0}^{V-1}F^{(0)}_{\text{i}}e^{-j2\pi\frac{mv}{V}}=F^{(0)}_{\text{R}}+jF^{(0)}_{\text{I}},$ (9)
where $F^{(0)}_{\text{o}}\in\mathbb{C}^{T\times D\times M}$, and $m\in[0,M-1]$ represents the frequency index, determining the frequency components extracted along the $V$ dimension. To process the complex-valued FFT outputs, we propose a spiking complex convolution module that preserves and utilizes amplitude and phase information, enhancing the network’s ability to capture detailed frequency-domain features:
$F^{(0)}_{\text{o},\text{R}}=\mathrm{BN}[\operatorname{Conv1d}(F^{(0)}_{\text{R}})],F^{(0)}_{\text{o},\text{I}}=\mathrm{BN}[\operatorname{Conv1d}(F^{(0)}_{\text{I}})],$ (10)
$F^{(0)}_{\text{o},k}=\text{SN}\left(F^{(0)}_{\text{o},\text{R}}\pm_{k}F^{(0)}_{\text{o},\text{I}}\right),\quad k\in\{1,2\}.$ (11)
where ±$\pm_{k}$ represents addition ($+$) when $k=1$, and subtraction (--) when $k=2$. The fused output is then computed as:
$F^{(0)}=\Phi\left(F^{(0)}_{\text{o},1}\right)+\Phi\left(F^{(0)}_{\text{o},2}\right).$ (12)
The output $F^{(0)}$ represents the extracted frequency-domain features, serving as an essential input for the subsequent layers of the network.
A complete 1D-SGC-FSC process can then be expressed as:
$X^{(l)}=F^{(l)}+G^{(l)}+\operatorname{res}\left(G^{(l-1)}\right),$ (13)
where, $X^{(l)}$ denotes the output of the $l$-th layer, combining the frequency-domain features $F^{(l)}$, temporal-domain features $G^{(l)}$, and a residual connection from the previous layer $\operatorname{res}\left(G^{(l-1)}\right)$.
III-B Multi-Scale Wavelet Transform Feature Fusion module
We formulate the MWTF module to capture the temporal dynamics and frequency-specific characteristics of spiking-form features, with its outputs fused with backbone outputs to enhance classification performance, as illustrated in Figure 1(c).
We first construct filters $\Lambda_{0}$, $\Lambda_{1}$, $\Gamma_{0}$, and $\Gamma_{1}$ based on Legendre polynomials $P_{k}(x)$ (details are provided in Appendix V-A) to decompose spiking-form features. The parameter $k$ specifies the number of frequency coefficients extracted, influencing the spectral resolution. We apply a linear (LN) transformation to the spiking features from the final layer $L$ of the backbone network to align the output tensor with the $k$ coefficients, ensuring compatibility with the filters.
$X_{\text{down}}=\operatorname{LN}\left(SN\left(X^{(L)}\right)\right)\in\mathbb{R}^{T\times k\times V}.$ (14)
To ensure compatibility with the wavelet transform, we verify whether the temporal dimension $T$ is a power of two. If not, $X_{\text{down}}$ is padded to the nearest power of two:
$X_{\text{extra}}=X_{\text{down}}[0:(2^{\lceil\log_{2}(T)\rceil}-T),:,:],$ (15)
$X_{\text{extended }}=\operatorname{concat}(X_{\text{down}},X_{\text{extra}},\text{ axis }=1),$ (16)
The extended spike-form feature $X_{\text{extended }}$ is decomposed by the filter matrices $\Lambda_{0}$, $\Lambda_{1}$, $\Gamma_{0}$, and $\Gamma_{1}$ into detail coefficients $D_{j}$, capturing high-frequency detailed features, and scaling coefficients $S_{j}$, retaining low-frequency trend features. The Algorithm 2 in Appendix V-B2 details the complete process.
To enhance the interaction between detail coefficients $D_{j}$ and scaling coefficients $S_{j}$ across decomposition levels $j$, we redefine the spiking cross-attention operation as:
$D_{j}^{\text{out}}=\left(\Phi_{Q}(D_{j})\cdot\Phi_{K}(D_{j})^{T}\right)\Phi_{V}(S_{j})*s,$ (17)
$S_{j}^{\text{out}}=\left(\Phi_{Q}(S_{j})\cdot\Phi_{K}(S_{j})^{T}\right)\Phi_{V}(D_{j})*s,$ (18)
where the temporal resolution is halved at each level. At the final level $J$, the scaling coefficient $S_{J}$ bypasses cross-attention and is directly upsampled to match the temporal resolution $T/2$:
$\hat{S}_{J}=\text{Upsample}(S_{J},T/2).$ (19)
All previous levels’ outputs $\{D_{j}^{\text{out}},S_{j}^{\text{out}}\}_{j=0}^{J-1}$ are similarly upsampled, and the aggregated feature map is computed as:
$X_{\text{MLO}}=\operatorname{GAP}\left(\Phi\left(\sum_{j=0}^{J-1}\hat{D}_{j}+\sum_{j=0}^{J}\hat{S}_{j}\right)\right).$ (20)
The final feature map $X_{\text{MLO}}$ integrates dynamic spiking features and spectral characteristics, which are fused with the output $X^{(L)}$ from the last backbone layer undergoing $SN$ and GAP. Then the fused feature proceed to a fully connected (FC) layer for classification:
$\hat{y}=\operatorname{FC}\left(\operatorname{GAP}\left(SN\left(X^{(L)}\right)\right)+\beta X_{\text{MLO}}\right),$ (21)
where $\beta$ is a learnable parameter. Finally, the cross-entropy loss is employed to optimize the classification task.
IV Experiments
We evaluate the Signal-SGN model on three skeleton-based action recognition datasets: NTU-RGB+D [30], NTU-RGB+D 120 [31], and NW-UCLA [32]. Detailed descriptions of the datasets, along with hyperparameter design, FLOPs, synaptic operations (SOPs), and theoretical energy consumption calculations, are provided in the Appendix V-C, V-D and V-E.
IV-A Comparison with the State-of-the-Art
TABLE I: Comparative results on NTU RGB+D and NTU-RGB+D 120. We evaluate our model in terms of classification accuracy (%). Xs and Xv represent cross-subject and cross-view splits.
Model (ANN/SNN) Param. (M) FLOPs (G) SOPs (G) Power (mJ) Xs (60) Xv (60) Xs (120) Xv (120)
Part-aware LSTM [33] - - - - 62.9 70.3 25.5 26.3
Spatio-Temporal LSTM [34] - - - - 61.7 75.50 55.7 57.9
ST-GCN [6] 3.10 3.48 - 16.01 81.5 88.3 70.7 73.2
2S-AGCN [7] 3.48 37.32 - 182.87 88.5 95.1 82.5 84.3
Shift-GCN [5] - 10.00 - 46.0 90.7 96.5 85.3 86.6
MS-G3D [35] 3.19 48.88 - 239.51 91.5 96.2 86.9 88.4
CTR-GCN [9] 1.46 7.88 - 36.25 92.4 96.8 88.9 90.6
Spikformer [13] ICLR 2023 4.78 24.07 1.69 2.17 73.9 80.1 61.7 63.7
Spike-driven Transformer [14] NIPS 2023 4.77 23.50 1.57 1.93 73.4 80.6 62.3 64.1
Spike-driven Transformer V2 [15] ICLR 2024 11.47 38.28 2.59 2.91 77.4 83.6 64.3 65.9
Spiking Wavelet Transformer [29] ECCV 2025 3.24 19.3 1.48 2.01 74.7 81.2 63.5 64.7
MK-SGN [17] 2.17 7.84 0.68 0.614 78.5 85.6 67.8 69.5
Signal-SGN (Joint) 1.74 1.62 0.314 0.372 80.5 87.7 69.2 72.1
Signal-SGN (Bone+Joint) 1.74 3.24 0.628 0.744 82.5 89.2 71.3 74.2
Signal-SGN (4 ensemble) 1.74 6.48 1.28 1.488 86.1 93.1 75.3 77.9
The Param. represents the model size for the network itself. For ANN, the metrics such as FLOPs, power, and accuracy are reported and calculated based on the maximum ensemble of modalities as specified in the referenced paper.
TABLE II: Comparative results on NW-UCLA. We evaluate our model in terms of classification accuracy (%).
Model Param. FLOPs SOPs Power Acc
(ANN/SNN) (M) (G) (G) (mJ) (%)
Ensemble TS-LSTM [3] - - - - 89.2
DC-GCN+ADG [18] 2.48 3.6 - 16.56 95.3
CTR-GCN [9] 1.46 2.32 - 10.67 96.5
Spikformer [13]ICLR 2023 4.78 8.55 0.513 0.673 85.4
Spike-driven Transformer [14]NIPS 2023 4.77 8.22 0.501 0.643 83.4
Spike-driven Transformer V2 [15]ICLR 2024 11.47 11.96 0.717 0.841 89.4
Spiking Wavelet Transformer [29]NIPS 2023 3.24 7.36 0.438 0.573 86.7
MK-SGN [17] 2.17 3.66 0.165 0.207 92.3
Signal-SGN (Joint) 1.74 0.67 0.161 0.147 92.7
Signal-SGN (Bone+Joint) 1.74 1.24 0.322 0.294 93.1
Signal-SGN (4 ensemble) 1.74 2.48 0.644 0.588 95.9
Several SOTA methods have adopted multi-stream fusion strategies [7, 5, 9]. For a fair comparison, we implement the same four-stream fusion approach for our Signal-SGN, incorporating ”joint”, ”bone,” ”joint motion”, and ”bone motion.” We compare Signal-SGN with state-of-the-art methods across three datasets. To compare with SOTA SNN models[13, 14, 15, 29], we perform unimodal reproductions using the same hierarchical architecture as MK-SGN[17], demonstrating the superiority of our method. The parameters of our Signal-SGN model, including FLOPs, SOPs, energy consumption, and accuracy for both single-modal and multi-modal scenarios, are presented in Tables I and II. In the NTU-RGB+D dataset, we resize the skeleton sequences to 16 frames, our model outperforms SOTA SNN models in accuracy while significantly reducing FLOPs, SOPs, and theoretical energy consumption when using single-modality joint data. Furthermore, compared to classical ANN models, our model achieves a higher accuracy than ST-GCN when integrating four modalities, which is quite rare for SNN. Our model also requires fewer FLOPs than the lightweight CTR-GCN and Shift-GCN models. In terms of theoretical energy consumption, our model reduces the energy usage to less than 5% compared to GCN-based methods, which is highly significant in practical applications.
In the NW-UCLA dataset, we also resize the skeleton sequences to 16 frames. Our model achieves an accuracy of 95.9%, which is only 0.6% lower than the classic CTR-GCN model while reducing theoretical energy consumption to 5.51% of CTR-GCN’s. This further demonstrates the effectiveness and energy-saving capabilities of our model. Additionally, removing the extra spike time step dimension reduces FLOPs and improves memory efficiency, making the model more trainable on standard devices compared to the previous SGN model.
IV-B Ablation Studies
To evaluate the contributions of the primary components in Signal-SGN, we conducted ablation studies on the NTU-RGB+D (cross-subject) dataset using joint data, with all skeleton sequences resized to 16 frames. Table III summarizes classification performance for different main module configurations. Appendix V-F presents a detailed analysis.
TABLE III: Comparison of experimental results under different module settings
Method 1D-SGC FSC MWTF Param. FLOPs/SOPs Acc
(M) (G) (%)
Ours - - 1.03 1.12/0.252 70.5
- 1.65 1.60/0.312 78.3
1.74 1.62/0.314 80.5
When only the 1D-SGC module is used, the model achieves an accuracy of 70.5%, demonstrating its ability to extract spiking-form graph features while maintaining low parameter count (1.03M) and computational costs (1.12G FLOPs/0.252G SOPs). Introducing FSC module significantly improves accuracy to 78.3%, highlighting the role of frequency-domain feature extraction in enhancing representations. The addition results in a moderate increase in parameter count to 1.65M and computational costs to 1.60G FLOPs/0.312G SOPs. Finally, incorporating the MWTF module further boosts accuracy to 80.5%. The improvement comes with a marginal parameter increase (1.74M) and computational costs (1.62G FLOPs/0.314G SOPs), demonstrating the MWTF module’s effectiveness in capturing multi-scale frequency features and fusing temporal and spectral information. These results validate the complementary roles of the 1D-SGC, FSC, and MWTF modules. The 1D-SGC module focuses on temporal-domain graph features, while the FSC module captures frequency-domain information. The MWTF module further enhances the model by integrating multi-scale frequency features, establishing Signal-SGN as an efficient framework for skeleton-based action recognition.
V Conclusion
We propose a novel Signal-SGN, a SNN-based framework designed to learn dynamic features in both the temporal and frequency domains for skeleton-based action recognition. The model comprises multiple stacked 1D-SGC-FSC layers, which utilize SGN to extract temporal-domain features from skeletons converted into spike forms and employ FFT and spiking complex convolution to extract frequency-domain feature. Additionally, the MWTF module, decomposes the spiking-form feature into distinct frequency components for robust feature fusion, enhancing classification performance, significantly. The final model surpasses SOTA SNNs in various aspects and achieves accuracy comparable to traditional ANN models while maintaining extremely low energy consumption, demonstrating its effectiveness and efficiency in skeleton-based action recognition.
Appendix
In the appendix, we first provide the supplemental detail description of Legendre Polynomial-Based Filters and pseudocode of the proposed methods, followed by a detailed introduction to the dataset. Then, we describe the hyperparameter settings and training details.
V-A Legendre Polynomial-Based Filters
Legendre polynomials $P_{k}(x)$, defined over the interval $[-1,1]$, provide a robust orthogonal basis for constructing frequency-domain filters aimed at decomposing spike-like features in temporal signals. Their orthogonality condition ensures that different polynomial orders yield mutually independent basis functions. The polynomials are generated recursively using the relation:
$P_{0}(x)=1,\quad P_{1}(x)=x,$ (22)
$(k+1)P_{k+1}(x)$ $=(2k+1)xP_{k}(x)-kP_{k-1}(x).$
Their orthogonality condition,as:
[…]
This property is particularly advantageous for designing a set of filters that can isolate specific frequency components with minimal interference.
To apply these polynomials to time-domain signals sampled over $[0,T]$, we first normalize the time index:
$u=\frac{t}{T},\quad u\in[0,1].$ (24)
We then map $[0,1]$ to $[-1,1]$:
$x=2u-1.$ (25)
This transformation aligns the original time domain with the canonical domain of the Legendre polynomials, ensuring straightforward evaluation of $P_{k}(x)$ at the corresponding sampling points.
Within this framework, low-frequency (“low-pass”) filters are derived from lower-order Legendre polynomials, which vary slowly across $[-1,1]$. To ensure proper normalization, each polynomial is scaled by $\sqrt{2k+1}$. Introducing a half-period shift $\Delta=0.5$ in the normalized time domain allows us to construct two distinct low-pass filters by evaluating the polynomials at slightly different points. Specifically, the low-pass filters are defined as:
$\Lambda_{0}[k,t]=\sqrt{2k+1}\,P_{k}(2u-1),$ (26)
$\Lambda_{1}[k,t]=\sqrt{2k+1}\,P_{k}(2(u+0.5)-1)=\sqrt{2k+1}\,P_{k}(2u).$ (27)
Here, $\Lambda_{0}[k,t]$ and $\Lambda_{1}[k,t]$ share the same polynomial family but are evaluated at $12u-$ 1 and $2u$, respectively, introducing a phase-like shift that distinguishes the two filters in their temporal and spectral responses.
For high-frequency (“high-pass”) filters, the ideal construction involves deriving them from multi-resolution analysis (MRA) conditions to ensure orthogonality and perfect reconstruction. Nonetheless, a representative scheme, consistent with the normalization above, is given by:
$\Gamma_{0}[k,t]=\sqrt{2}\,\Lambda_{0}[k,t],$ (28)
$\Gamma_{1}[k,t]=\sqrt{2}\,\Lambda_{1}[k,t].$ (29)
Taken together, these four filters $(\Lambda_{0},\Lambda_{1},\Gamma_{0},\Gamma_{1})$ can decompose the signal into distinct frequency bands, enabling a nuanced analysis of spike-like features. By carefully mapping temporal samples into the Legendre polynomial domain, employing a normalized and shifted set of low-pass filters, and incorporating a high-pass counterpart, this approach provides a stable, orthogonal, and frequency-selective analytical framework. This framework facilitates effective characterization, improved interpretability, and subsequent analyses of time-varying signals in a multi-resolution context.
V-B Algorithm
This section presents the essential pseudocode and detail description underpinning the computational framework discussed in this work.
V-B1 Spiking Neuron Dynamics Simulation
The Spiking Neuron Dynamics describes the core mechanisms governing the temporal dynamics of spiking neurons within the network, outlined in Algorithm 1.
Algorithm 1 Spiking Neuron Dynamics Simulation
1:  Input: $G\in\mathbb{R}^{T\times D\times V}$
2:  Parameters: $\tau$, $V_{\text{rest}}$, $R$, $V_{\text{th}}$
3:  Output: $G_{o}\in\mathbb{R}^{T\times C\times V}$
4:  Initialize $V(t)$ and $G_{o}$
5:  for $t=0$ to $T-1$ do
6:     $G_{i}(t)\leftarrow G[t,:,:]$
7:     $\tau\frac{dV(t)}{dt}\leftarrow-\left(V(t)-V_{\text{rest}}\right)+R\cdot G_{i}(t)$
8:     if $V(t)\geq V_{\text{th}}$ then
9:        $G_{o}[t,:,:]\leftarrow 1$
10:        $V(t)\leftarrow V_{\text{rest}}$Reset voltage after spike}
11:     else
12:        $G_{o}[t,:,:]\leftarrow 0$
13:     end if
14:  end for
15:  return $G_{o}\in\mathbb{R}^{T\times D\times V}$
The input, $G\in\mathbb{R}^{T\times D\times V}$, represents the temporal sequence of feature maps for $T$ time steps, where $D$ and $V$ denote the feature dimension and the number of vertices (e.g., joints in skeleton data), respectively. The algorithm computes the output $G_{o}\in\mathbb{R}^{T\times C\times V}$, which represents the spike emissions over time, with $C$ corresponding to the output channel dimension.
At each time step $t$, the membrane potential $V(t)$ is updated based on the input current $G_{i}(t)$ and the parameters governing spiking neuron dynamics, including the membrane time constant $\tau$, the resting potential $V_{\text{rest}}$, the resistance $R$, and the firing threshold $V_{\text{th}}$. Specifically, the potential evolves according to the equation:
$\tau\frac{dV(t)}{dt}=-\left(V(t)-V_{\text{rest}}\right)+R\cdot G_{i}(t),$ (30)
where $G_{i}(t)=G[t,:,:]$ is the input tensor for time step $t$. If the membrane potential $V(t)$ exceeds the threshold $V_{\text{th}}$, a spike is emitted ($G_{o}[t,:,:]=1$), and the potential is reset to $V_{\text{rest}}$. Otherwise, no spike is generated ($G_{o}[t,:,:]=0$).
This mechanism enables the simulation of spike-based information transmission, where $G_{o}$ captures the spiking behavior at each time step. The tensor $G_{o}$ has the same temporal and spatial resolution as the input but reflects the binary spike activations, effectively encoding temporal dynamics into discrete spike events.
The iterative process ensures that spiking neurons dynamically respond to varying inputs while adhering to biologically inspired principles. Subsequent layers of spiking neurons follow similar dynamics, allowing the network to propagate spike-based representations through deeper processing stages.
V-B2 Iterative Decomposition
The Iterative Decomposition algorithm describes the process of decomposing spike-form features into their frequency domain representations. This hierarchical decomposition enables the separation of temporal dynamics into high-frequency detail coefficients and low-frequency scaling coefficients, which are critical for efficient feature extraction and multi-resolution analysis in SNNs.
Algorithm 2 Iterative Decomposition
1:  $X_{0}\leftarrow X_{\text{extended}}$
2:  for $j=0$ to $J-1$ do
3:     $X_{\text{even}}^{(j)}\leftarrow X_{j}[::2,:,:]$
4:     $X_{\text{odd}}^{(j)}\leftarrow X_{j}[1::2,:,:]$
5:     $D_{j}\leftarrow\Gamma_{0}\cdot X_{\text{even}}^{(j)}+\Gamma_{1}\cdot X_{\text{odd}}^{(j)}$
6:     $S_{j}\leftarrow\Lambda_{0}\cdot X_{\text{even}}^{(j)}+\Lambda_{1}\cdot X_{\text{odd}}^{(j)}$
7:     $X_{j+1}\leftarrow S_{j}$
8:  end for
9:  $S_{J}\leftarrow X_{J}$
10:  return $\{D_{j}\}_{j=0}^{J-1}$ and $\{S_{j}\}_{j=0}^{J}$
The iterative decomposition process begins with the extended spike-form feature tensor $X_{\text{extended}}$, which serves as the initial input. The tensor is initialized as $X_{0}$, providing the starting representation for decomposition across $J$ levels.
At each decomposition level $j$, the input tensor $X_{j}$ is divided along its temporal dimension into even and odd indexed components, denoted as $X_{\text{even}}^{(j)}$ and $X_{\text{odd}}^{(j)}$, respectively. These components are independently processed to compute the detail coefficients $D_{j}$ and scaling coefficients $S_{j}$ as follows:
$D_{j}=\Gamma_{0}\cdot X_{\text{even}}^{(j)}+\Gamma_{1}\cdot X_{\text{odd}}^{(j)},$ (31)
$S_{j}=\Lambda_{0}\cdot X_{\text{even}}^{(j)}+\Lambda_{1}\cdot X_{\text{odd}}^{(j)}.$ (32)
where, $\Gamma_{0},\Gamma_{1}$ are high-pass filter matrices used to extract the detail coefficients $D_{j}$, which capture high-frequency temporal variations. Conversely, $\Lambda_{0},\Lambda_{1}$ are low-pass filter matrices that compute the scaling coefficients $S_{j}$, preserving the low-frequency temporal information.
After computing $D_{j}$, the scaling coefficients $S_{j}$ are propagated to the next level by setting $X_{j+1}=S_{j}$. This iterative decomposition continues across all $J$ levels, refining the temporal representation at each step. Upon completing the $J$-th level, the final scaling coefficients $S_{J}$ are obtained. The process outputs the set of detail coefficients $\{D_{j}\}_{j=0}^{J-1}$ and scaling coefficients $\{S_{j}\}_{j=0}^{J}$, providing a hierarchical representation of the temporal dynamics.
Upon completion, the algorithm outputs a set of detail coefficients $\{D_{j}\}_{j=0}^{J-1}$ and the final scaling coefficients $S_{J}$, collectively encapsulating the hierarchical temporal structure of the input. The complete iterative decomposition process is outlined in Algorithm 2, providing a systematic and computationally efficient method for processing spike-form freature.
V-C Datasets
V-C1 NTU-RGB+D
NTU-RGB+D[30] is a comprehensive dataset widely used in the field of skeleton-based action recognition. It contains 60 action classes and 56,000 video clips, making it a robust benchmark for evaluating action recognition models. The dataset provides detailed 3D skeleton data, RGB frames, depth maps, and infrared sequences, covering a diverse range of daily activities such as eating, walking, and interacting with objects. The skeleton data consists of 25 joints per subject, captured using the Microsoft Kinect v2 sensor, which allows for precise modeling of human movement. This dataset is particularly valuable for researchers because it includes a variety of subjects performing actions from multiple viewpoints, offering a rich source of information to develop and test models capable of understanding complex human behaviors.
V-C2 NTU-RGB+D 120
NTU-RGB+D 120[31] is an extended version of the original NTU-RGB+D dataset, expanding the number of action classes from 60 to 120 and increasing the total number of video clips to over 114,000. This extension includes additional action categories that encompass more complex and subtle human interactions, such as ”taking a selfie” and ”writing on a board.” With 106 distinct subjects and more diverse camera angles, NTU-RGB+D 120 offers a more challenging and comprehensive benchmark for evaluating the scalability and generalization capabilities of action recognition models. The dataset retains the detailed skeleton tracking of its predecessor, providing 25 joints per person and ensuring that the data remains highly informative for developing models that require an in-depth understanding of spatial-temporal dynamics in human actions. As one of the largest skeleton-based action recognition datasets available, NTU-RGB+D 120 is instrumental in pushing the boundaries of research in this field.
V-C3 NW-UCLA
The NW-UCLA dataset[32] is a smaller but uniquely challenging dataset consisting of 10 action classes and 1,494 video clips. It is specifically designed to test the robustness of action recognition models against viewpoint variations. The dataset features actions like jumping, waving, and running, captured from multiple viewpoints using three Kinect cameras placed at different angles. Each action is performed by 10 subjects, offering diverse skeleton data with 20 joints per person. This setup mimics real-world scenarios where actions are observed from different perspectives, making it a critical dataset for evaluating how well models can maintain performance when faced with varying observational conditions. The NW-UCLA dataset is ideal for researchers aiming to enhance the adaptability and resilience of their models to changes in camera viewpoints, which is essential for applications in surveillance, human-computer interaction, and robotics.
V-D Model Implementation and Configuration
The models used for conducting experiments are implemented using PyTorch [^1] https://github.com/rwightman/pytorch-image-models, a classical deep learning framework known for its flexibility and efficiency, and SpikingJelly [^2] https://github.com/fangwei123456/spikingjelly, a specialized library for spiking neural network research. Our implementation leverages the powerful features of these tools to effectively model and simulate the complex dynamics of spiking neural networks. The training process is conducted on four NVIDIA V100 GPUs, which provide the necessary computational power and memory to handle the large datasets and intricate computations involved in our experiments. This setup ensures that our models are trained efficiently, allowing us to explore the full potential of our approach in a scalable and robust manner.
TABLE IV: The parameter Settings of the Backbone
Level of Layers Channel Size
Input Output
1 3 64
2 64 64
3 64 128
4 128 256
In this section, we also give the specific hyperparameters of the model and training settings in all experiments. Table IV. presents the parameter settings of the backbone network, detailing the input and output channel sizes across different layers to illustrate the model’s architectural configuration.
The various hyperparameter settings used in our experiments are also shown in the TableV.
TABLE V: Hyper-parameter training settings of Signal-SGN.
Parameter NTU-RGB+D NTU-RGB+D NW-UCLA
60 120
Learning Rate 0.1/0.1 0.1/0.1 0.1/0.1
Decay rate 0.1/0.1 0.1/0.1 0.1/0.1
Batch Size 64 64 128
Training Epochs 90 90 90
Step 50 50 50
Dropout Rate 0.3 0.3 0.3
Weight Decay 1e-4 1e-4 1e-4
$\omega$ 0.03 0.03 0.03
Optimizer SGD SGD SGD
V-E Power Consumption
We assume that multiply-and-accumulate (MAC) and accumulate (AC) operations are implemented on the 45 nm technology node with $E_{MAC}$ = 4.6 pJ and $E_{AC}$ = 0.9 pJ. The $\mathrm{FLOPs}$ for Signal-SGN and all reproduced models are calculated using FVCore [^3] https://github.com/facebookresearch/fvcore, providing accurate estimates for each model component. These FLOPs are then combined with the firing rate and spike time steps to compute SOPs as:
$\mathrm{SOPs}(l)=fr\times T\times\mathrm{FLOPs}(l),$ (33)
where $fr$ is firing rate, $T$ is the spike time steps, $\operatorname{FLOPs}(l)$ refers to the floating-point operations of $l$, defined as the number of MAC operations performed in $l$ block or layer, and $\operatorname{SOPs}$ are the number of spike-based AC operations.
The model’s theoretical energy consumption is described as:
$\begin{gathered}E_{\text{Signal-SGN }}=E_{EAC}\times\mathrm{FL}_{\mathrm{SNN}\text{ Conv }}^{1}+\\ E_{AC}\times\left[\left(\sum_{n=2}^{N}\mathrm{SOP}_{\mathrm{SNN}\text{ Conv }}^{n}+\sum_{m=1}^{M}\mathrm{SOP}_{\mathrm{SNN}\text{ FC }}^{m}\right.\right.\\ \left.\left.+\sum_{l=1}^{L}\mathrm{SOP}_{\mathrm{SNN}\text{ SSA }}^{l}+\sum_{f=1}^{F}\mathrm{SOP}_{\mathrm{SNN}\text{ FFT }}^{f}\right)+\mathrm{SOP}_{\mathrm{SNN}\text{ MWTF }}\right]\end{gathered}$ (34)
where $\mathrm{FL}_{\mathrm{SNN}\text{ Conv }}^{1}$ is the first layer to encode skeleton into spike-form. Then the SOPs of $N$ SNN Conv layers, $M$ SNN Fully Connected Layer (FC), $L$ SSA , $F$ SSA and MWTF are added together are added together and multiplied by $E_{AC}$. The theoretical energy consumption for ANN and SNN models is calculated as follows:
For ANNs, the energy consumption of a block $b$ is given by:
$\text{Power}(b)=4.6\,\mathrm{pJ}\times\mathrm{FLOPs}(b).$ (35)
For SNNs, the energy consumption of a block $b$ is given by:
$\text{Power}(b)=0.9\,\mathrm{pJ}\times\mathrm{SOPs}(b).$ (36)
V-F More detail of Ablation Studies
In this section, we provide additional details regarding the ablation studies conducted in our research. These experiments are designed to investigate the contribution of various components within our framework and to demonstrate the effectiveness of specific design choices. By systematically removing or altering key elements, we aim to provide a deeper understanding of their impact on the overall performance.
3D coordinate normalization
3D skeleton data often spans multiple quadrants and includes negative values, as shown in Figure 2, creating challenges for SNNs, which are inherently designed for non-negative numerical inputs. Negative coordinates lead to significant information loss and degrade model performance. To address this limitation, 3D coordinate normalization is applied to map skeleton data into a fixed and bounded range, ensuring numerical consistency and preserving essential information. Normalization facilitates stable data propagation through the network by mitigating discrepancies arising from varying magnitudes and signs.
Figure 2: 3D Skeleton Visualization
The analysis focuses primarily on accuracy as the evaluation metric, as changes in parameters and FLOPs introduced by normalization are minimal. By isolating the impact of 3D coordinate normalization on accuracy, the study demonstrates that normalization significantly enhances model performance without introducing substantial computational overhead. Controlled experiments under various module configurations reveal consistent accuracy improvements, underscoring this approach’s effectiveness in processing complex skeleton data. Detailed results presented in the accompanying table highlight the practical benefits of normalization in achieving higher accuracy with negligible computational trade-offs.
TABLE VI: Impact of 3D Coordinate Normalization on Accuracy
3D Coord. Norm. 1D-SGC FSC MWTF Acc (%)
- - - 69.7
- - 77.0
- 79.3
- - 70.5
- 78.3
80.5
Figure 3: Comparison of confusion matrices for full-module integration setup with and without 3D coordinate normalization
The results in Table 3 demonstrate the significant role of 3D coordinate normalization in improving accuracy across all module configurations. For the baseline configuration (1D-SGC only), normalization increases accuracy by 0.8%, while in more complex setups with FSC and MWTF modules, the improvements reach 1.3% and 1.2%, respectively. This indicates that normalization effectively stabilizes the input data by addressing inconsistencies in coordinate ranges, enabling the model to better utilize its capacity. Importantly, the improvements in accuracy are achieved without notable changes in parameters or computational cost, highlighting normalization as a practical and efficient enhancement for processing 3D skeleton data. To further analyze the impact of 3D coordinate normalization, we selected the full-module integration setup (MWTF enabled) from the third and sixth rows of Table VI and visualized the corresponding confusion matrices.
Different Settings for MWTF
To investigate the impact of different configurations of $k$ (frequency resolution) and $J$ (decomposition levels) on the performance of the MWTF module within Signal-SGN, we conduct a series of controlled experiments. The overall architecture of Signal-SGN is kept unchanged, ensuring that any observed variations in performance can be attributed solely to changes in the MWTF parameters. By systematically adjusting $k$ and $J$, we aim to analyze how the trade-off between frequency granularity and decomposition depth affects the network’s ability to extract and fuse temporal features effectively.
TABLE VII: Comparison of experimental results under different MWTF settings ($k$ and $J$)
MWTF $k$ $J$ Param. (M) FLOPs/SOPs (G) Acc (%)
- - - 1.65 1.60/0.312 78.3
4 1 1.68 1.61/0.313 78.9
8 1 1.73 1.62/0.314 79.8
8 2 1.74 1.62/0.315 80.3
8 3 1.74 1.62/0.317 80.5
16 3 1.83 1.64/0.320 80.5
The experimental results in Table VII highlight the impact of varying $k$ (frequency resolution) and $J$ (decomposition levels) on the performance of the MWTF module within Signal-SGN. With the overall architecture unchanged, the table isolates the influence of MWTF parameters, providing insight into their effects on accuracy, computational complexity, and parameter efficiency.
When MWTF is disabled, the model achieves a baseline accuracy of 78.3% with a parameter count of 1.65M and computational costs of 1.60/0.312 G (FLOPs/SOPs). Enabling MWTF with different $k$ and $J$ configurations demonstrates significant improvements. As $k$ increases from $4$ to $8$, accuracy improves substantially, with the model achieving 79.8% ($k=8,J=1$). Further increasing $k$ to $16$ yields diminishing returns, with accuracy peaking at 80.5% for $k=16,J=3$.
Similarly, increasing $J$ allows for deeper decomposition of temporal features, which enhances performance up to a point. For example, with $k=8$, accuracy improves from 79.8% ($J=1$) to 80.7% ($J=3$). This configuration achieves the best accuracy while maintaining a parameter count of 1.74M and computational cost of 1.62/0.317 G. Further increasing $k$ to $16$ ($J=3$) slightly reduces accuracy to 80.5%, indicating that excessively high frequency resolution or decomposition depth might lead to redundant features or overfitting.
Overall, the results demonstrate that the MWTF module significantly enhances the performance of Signal-SGN when configured optimally. The configuration $k=8,J=3$ emerges as the most effective setting, balancing frequency granularity, decomposition depth, and computational efficiency, and achieving the best accuracy of 80.7%. This underscores the importance of selecting appropriate MWTF parameters to maximize feature extraction and model performance.

### v5

@@
 Signal-SGN: A Spiking Graph Convolutional Network for Skeleton Action Recognition via Learning Temporal-Frequency Dynamics
-Abstract
-For skeleton-based action recognition, Graph Convolutional Networks (GCNs) are effective models.
+Abstract.
+For multimodal skeleton-based action recognition, Graph Convolutional Networks (GCNs) are effective models.
 Still, their reliance on floating-point computations leads to high energy consumption, limiting their applicability in battery-powered devices.
@@
 Experiments across three large-scale datasets reveal Signal-SGN exceeding state-of-the-art SNN-based methods in accuracy and computational efficiency while attaining comparable performance with GCN methods and significantly reducing theoretical energy consumption.
-Index Terms:
 Skeleton-based Action Recognition, Spiking Neural Networks, Frequency Domain Transform
-I Introduction
-Skeleton-based action recognition, as a critical subfield of human action recognition, has made significant strides in recent years.
-Leveraging sparse representations obtained from pose estimation and sensor-based data, it enables effective classification of human actions [1].
-Numerous methods have been proposed for this task, including CNNs [2], LSTMs [3], Transformers [4], and Graph Convolutional Networks (GCNs) [5].
-Among them, ST-GCN is a pioneering method that laid the foundation for GCN-based approaches [6].
-Several methods,such as 2s-AGCN [7] and Dynamic GCN [8], have been proposed subsequently, further improving accuracy in action recognition.
-Lightweight models like Shift-GCN [5] and CTR-GCN [9] have further enhanced accuracy while reducing computational costs.
+ccs: Computing methodologies Activity recognition and understanding
+1. Introduction
+Multimodal skeleton-based action recognition aims to identify human actions by analyzing motion patterns from various joint-related modalities such as joints, bones, and their temporal displacements (Ren et al., 2024).
+To tackle this task, numerous methods have been proposed, including CNNs (Du et al., 2015; Caetano et al., 2019b; Xu et al., 2018; Li et al., 2019; Caetano et al., 2019a), LSTMs (Gao et al., 2022; Li et al., 2021, 2018), Transformers (Plizzari et al., 2021; Liu et al., 2023; Chen et al., 2022; Bai et al., 2022), and Graph Convolutional Networks (GCNs) (Yan et al., 2018; Shi et al., 2019; Cheng et al., 2020b; Chen et al., 2021; Wang et al., 2022).
+Among them, ST-GCN is a pioneering method that laid the foundation for GCN-based approaches (Yan et al., 2018).
+Subsequent methods, such as 2s-AGCN (Shi et al., 2019) and Dynamic GCN (Ye et al., 2020), have been proposed to achieve higher accuracy in action recognition.
+Lightweight models like Shift-GCN (Cheng et al., 2020b) and CTR-GCN (Chen et al., 2021) have optimized accuracy while significantly reducing computational costs.
 However, these methods are all based on artificial neural networks (ANNs), which rely on floating-point operations (FLOPs), leading to significant energy consumption that limits their practicality in energy-constrained applications.
+Figure 1.
+Motivating perspective of Signal-SGN.
+Skeleton data, once encoded by the spiking neural network, becomes a sparse binary representation interpretable as a multi-dimensional discrete signal.
+Signal-SGN leverages this perspective to apply time-frequency analysis, revealing structured spectral-temporal patterns critical for action recognition.
 Spiking Neural Networks (SNNs), regarded as third-generation neural networks, utilize spike-form features encoded as 0s and 1s to mimic neural dynamics.
-Unlike ANNs, SNNs process information in an event-driven manner, enabling significantly lower energy consumption while maintaining efficient computation [10].
-Various SNN models have been proposed, leveraging their energy efficiency, and have been applied to a wide range of tasks, including image classification and point cloud estimation [11, 12, 13, 14, 15].
+Unlike ANNs, SNNs process information in an event-driven manner, enabling significantly lower energy consumption while maintaining efficient computation (Ghosh-Dastidar S., 2009).
+Various SNN models have been proposed, leveraging their energy efficiency, and have been applied to a wide range of tasks, including image classification and point cloud estimation (Yamazaki et al., 2022; Tavanaei et al., 2019; Zhou et al., 2022; Yao et al., 2024b, a).
 Among these, Spiking Graph Networks (SGNs) combine graph convolution with SNNs, achieving remarkable energy savings.
-For instance, SGNs reduce energy consumption by nearly 99% compared to vanilla GCNs [16].
-Specifically, SGN has been utilized for skeleton action recognition, cutting energy consumption to less than 5% of GCNs while maintaining comparable accuracy [17].
-Despite its advantages, [17] treats dynamic skeleton data as static image data, artificially extending it along the spike time step dimension.
-This approach neglects the time-sensitive computational capabilities of SNNs, failing to effectively capture the inherent temporal dynamics of skeleton data.
+For instance, SGNs reduce energy consumption by nearly 99% compared to vanilla GCNs (Zhu et al., 2022).
+Specifically, SGN has been utilized for skeleton action recognition, cutting energy consumption to less than 5% of GCNs while maintaining comparable accuracy (Zheng et al., 2024).
+Despite its advantages, (Zheng et al., 2024) treats dynamic skeleton data as static image data, artificially extending it along the spike time step dimension.
+This approach neglects the temporal-sensitive computational capabilities of SNNs, failing to effectively capture the inherent temporal dynamics of skeleton data.
 As a result, it increases computational complexity, hampers performance, and poses significant challenges for efficient training on standard deep-learning hardware.
-Furthermore, SGNs, like other SNN-based models [13, 14, 15], perform a simplistic Global Average Pooling (GAP) over the spike time step dimension in the task head, failing to thoroughly analyze the discrete and aperiodic nature of spike-form features.
-This oversimplification limits the model’s ability to capture temporal patterns, reducing its task performance.
-To overcome these challenges, we propose Signal-SGN for skeleton action recognition, utilizing the temporal dimension of skeleton sequences as the spiking time step and representing skeleton data as spiking-form features with multi-dimensional discrete stochastic properties.
+Furthermore, SGNs, like other SNN-based models (Zhou et al., 2022; Yao et al., 2024b, a), perform a simplistic Global Average Pooling (GAP) over the spike time step dimension in the task head.
+This naive averaging overlooks the discrete and temporal dependencies of spike-form features, impairing the model’s capacity to analyze temporal structures.
+To overcome these challenges, we propose Signal-SGN for skeleton action recognition, utilizing the temporal dimension of skeleton sequences as the spiking time step and representing skeleton data as spiking-form features with multi-dimensional discrete stochastic properties, as shown in Figure1.
 The architecture comprises two primary modules: a one-dimensional Spiking Graph Convolution (1D-SGC) module and a Frequency Spiking Convolution (FSC) module.
@@
 We formulate a Multi-Scale Wavelet Transform Feature Fusion (MWTF) module to address the limitations of simple GAP, which fails to effectively capture the temporal dynamics and frequency-specific characteristics of spiking-form features.
-The module exploits discrete wavelet transform (DWT) to decompose features into distinct frequency components and integrates a spiking cross-attention mechanism for robust feature fusion.
+The module leverages discrete wavelet transform (DWT) to decompose features into distinct frequency components and employs a spiking cross-attention mechanism to effectively fuse these components for robust feature representation.
 The proposed module significantly improves classification performance.
@@
   • We propose a novel Signal-SGN network comprising multi-layer 1D-SGC-FSC that captures temporal-frequency domain spatial features of single-frame skeleton graphs and leverages the time-sensitive computational properties of SNNs to effectively model sequence-level temporal dynamics.
-  • We propose the MWTF module, which achieves multi-scale frequency-domain decomposition and interaction between frequency-specific features to extract dynamic spiking features, capture frequency-specific characteristics, and improve classification performance.
+  • We propose the MWTF module, which performs multi-scale temporal-frequency decomposition and integrates frequency-specific interactions to extract the temporal-frequency characteristics of spike-form features and enhance classification performance.
   • Extensive experiments across three large-scale demonstrate that Signal-SGN surpasses State-of-the-Art (SOTA) SNN methods in terms of computational costs, and accuracy while achieving results comparable to GCN-based methods with significantly lower energy consumption.
-II Related Work
-II-A Skeleton-Based Action Recognition.
+2. Related Work
+2.1. Skeleton-Based Action Recognition.
 Deep learning methods have played a significant role in skeleton-based action recognition.
-Early approaches in this field achieved significant success using CNNs[2] and RNNs[3].
-Subsequent research has primarily focused on using Transformers[4] and GCNs[6, 7, 5, 18, 9].
+Early approaches in this field achieved significant success using CNNs(Du et al., 2015; Caetano et al., 2019b; Xu et al., 2018; Li et al., 2019; Caetano et al., 2019a) and RNNs(Gao et al., 2022, 2022; Li et al., 2021, 2018).
+Subsequent research has primarily focused on using Transformers(Plizzari et al., 2021; Liu et al., 2023; Chen et al., 2022; Bai et al., 2022) and GCNs(Yan et al., 2018; Shi et al., 2019; Cheng et al., 2020b, a; Chen et al., 2021).
 These numerous GCN-based methods typically involve nearly multi-layer stacks of Graph Convolutional Networks and Temporal Convolutional Networks.
@@
 Such reliance leads to substantial energy consumption and fundamentally misaligns with practical requirements for energy-efficient solutions.
-II-B Spiking Neural Networks.
-SNNs, inspired by biological neural networks, offer an energy-efficient alternative to traditional neural networks [10].
-They can be implemented through two primary approaches: ANN-to-SNN conversion [19] and direct training using algorithms like STDP [20] or surrogate gradients [21].
-The fundamental units, such as Leaky Integrate-and-Fire (LIF) neurons, are notable for their simplicity and biological plausibility [22].
+2.2. Spiking Neural Networks.
+SNNs, inspired by biological neural networks, offer an energy-efficient alternative to traditional neural networks (Ghosh-Dastidar S., 2009).
+They can be implemented through two primary approaches: ANN-to-SNN conversion (Ho and Chang, 2021) and direct training using algorithms like STDP (Caporale and Dan, 2008) or surrogate gradients (Neftci E.
+O., 2019).
+The fundamental units, such as Leaky Integrate-and-Fire (LIF) neurons, are notable for their simplicity and biological plausibility (Teeter et al., 2018).
 The differential equation describes the dynamics of the LIF neuron:
-$\tau_{m}\frac{dV(t)}{dt}=-V(t)+R\cdot I(t),$ (1)
+(1) $\tau_{m}\frac{dV(t)}{dt}=-V(t)+R\cdot I(t),$
 where $V(t)$ is the membrane potential, $\tau_{m}$ is the time constant indicating potential decay, $R$ is the membrane resistance, and $I(t)$ is the input current.
@@
 The LIF neuron model accurately replicates the spiking dynamics characteristic of neurons within SNNs.
-Inspired by ANN architectures, corresponding SNN models have been proposed and applied to tasks such as image classification, event-driven processing, and EEG analysis[23, 24, 25].
-Among these, various spiking-based Transformers have played significant roles[13, 14, 15].
-Meanwhile, [17] introduces Spiking Graph Networks (SGNs) for skeleton-based action recognition for the first time, achieving promising results.
-However, [17] treats the skeleton sequence data as static and artificially extends them along the spike time step dimension, which increases storage and computational demands.
+Figure 2.
+Overview of the proposed Signal-SGN framework for skeleton-based action recognition.
+The model processes skeleton sequences through four main stages: a 3D Coordinate Normalization (3D C.N.) to stabilize input representations, a 1D-Spiking Graph Convolution (1D-SGC) module to capture topological dependencies in the spiking domain, a Frequency Spiking Convolution (FSC) module that extracts frequency-aware features via FFT and complex convolution, and a Multi-Scale Wavelet Transform Feature Fusion (MWTF) module that performs multi-level decomposition and integration using discrete wavelet transform and spiking cross-attention.
+Final classification is conducted through global average pooling (GAP) and a fully connected (FC) layer.
+Inspired by ANN architectures, corresponding SNN models have been proposed and applied to tasks such as image classification, event-driven processing, and EEG analysis(Cao Y., 2015; Pedretti et al., 2020; Cai et al., 2024).
+Among these, various spiking-based Transformers have played significant roles(Zhou et al., 2022; Yao et al., 2024b, a).
+Meanwhile, (Zheng et al., 2024) introduces SGN for skeleton-based action recognition for the first time, achieving promising results.
+However, (Zheng et al., 2024) treats the skeleton sequence data as static and artificially extends them along the spike time step dimension, which increases storage and computational demands.
 Additionally, SGNs fail to capture the discrete and non-periodic nature of spike-form features and rely on simple GAP to reduce the spike time step dimension before the classification head, limiting classification performance and accuracy.
-II-C Frequency Domain Analysis.
-In neural network design, frequency domain analysis is crucial for improving performance and robustness.
-The FFT and Wavelet Transform (WT) enhance feature extraction within CNNs for applications such as image compression [26].
-FFT also plays an important role in sequential data analysis within RNNs, which is particularly beneficial for applications such as speech recognition [27].
-GCNs integrate Fourier transforms (FT) to analyze graph signals [28].
-Meanwhile, SNNs combined with WT enable effective spatial-frequency domain learning, achieving energy efficiency and high performance [29].
-In the context of SNNs, skeleton sequences are represented as spiking-form feature, resembling discrete multi-dimensional stochastic signals.
-Exploring their frequency domain characteristics has significant potential for the analysis of skeleton action recognition within SNNs.
-However, research in the area is currently insufficient.
-Figure 1: The overview of the proposed Signal-SGN method.
-III Method
-This section presents the overall architecture of Signal-SGN, provided in Figure 1.
-It begins with a description of the backbone network, including the 1D-SGC module and the FSC module, which constitute the core components of the model.
-Following this, the MWTF module is explained, emphasizing its role for enhancing performance.
-III-A The Backbone of Signal-SGN
-As illustrated in Figure 1, Signal-SGN processes three-dimensional skeleton data $X\in\mathbb{R}^{C\times T\times V}$, where $C$ represents spatial coordinates $(x,y,z)$, $T$ denotes the number of frames and spike time steps, and $V$ is the number of joints per frame.
-The spatial coordinates, which contain both positive and negative values, pose challenges to SNNs due to rate coding and binarization, leading to information loss.
-To address this, the coordinates are normalized to $[0,1]$ using the formula $a^{\prime}=(a-a_{\min})/(a_{\max}-a_{\min})$, where $a\in\{x,y,z\}$.
-After normalization, the skeleton data is transformed into $X_{\text{i}}\in\mathbb{R}^{T\times C^{\prime}\times V}$, where $C^{\prime}$ corresponds to the normalized three-dimensional coordinates.
+2.3. Frequency Domain Analysis.
+Frequency-domain analysis has become an increasingly important tool across various deep learning architectures for enhancing robustness and representational capacity(Yi et al., 2023).
+The Fourier Transform (FT) and Wavelet Transform (WT) have been extensively adopted in CNNs for image compression and reconstruction (Li et al., 2020), in RNNs for modeling sequential dynamics such as speech (Jalayer M., 2021), and in GCNs for processing graph signals (Yu and Qin, 2020).
+Recent work also demonstrates that adaptive frequency-based token mixing can replace costly attention mechanisms while maintaining global context modeling capabilities (Tatsunami and Taki, 2024).
+Moreover, in camouflaged object detection, frequency-spatial entanglement learning has proven effective in mitigating the limitations of local spatial features (Sun et al., 2024).
+In the context of Vision Transformers, frequency-guided representations have also shown to improve generalization and robustness by modulating attention with spectral priors (Patro et al., 2023).
+In the context of SNNs, skeleton sequences are encoded as discrete multi-dimensional spike signals—akin to stochastic binary signals.
+Exploring their frequency characteristics offers promising potential for interpretable analysis and enhanced recognition.
+However, frequency-domain analysis for SNN-based skeleton action recognition remains underexplored and merits further attention.
+3. Method
+This section details the architecture of the proposed Signal-SGN, as illustrated in Figure 2.
+Given an input skeleton sequence, the model first applies a 3D Coordinate Normalization (3D C.N.) to standardize joint coordinates and enhance spiking encoding consistency.
+Then, the 1D-SGC module captures spatial dependencies of joints through spiking-based graph convolution.
+To further exploit frequency-domain information, FSC module transforms features via FFT and applies complex convolution for frequency-aware representation learning.
+Subsequently, the MWTF module decomposes spike-form features into multi-resolution frequency components using discrete wavelet transform, followed by cross-attentive fusion.
+The final prediction is made through GAP and a fully connected(FC) layer.
+3.1. 3D C.N.
+As shown in Figure 3, 3D skeleton data spans multiple spatial quadrants and includes both positive and negative coordinate values.
+This characteristic introduces challenges for SNNs, which are inherently designed for non-negative numerical inputs.
+In particular, the presence of negative values can disrupt rate-based spike encoding and lead to severe information loss due to binarization effects.
+To address this, we apply a per-channel normalization that maps each spatial coordinate to the range $[0,1]$.
+Given a skeleton sequence $X\in\mathbb{R}^{C\times T\times V}$, where $C=3$ denotes the spatial dimensions $(x,y,z)$, $T$ is the number of frames, and $V$ is the number of joints, we normalize each coordinate channel $c\in\{x,y,z\}$ independently across the full sequence utilizing:
+(2) $c^{\prime}=\frac{c-\min(c)}{\max(c)-\min(c)}$
+where $\min(c)$ and $\max(c)$ are computed over all joints and time steps for each coordinate channel.
+This process preserves the relative geometric structure while ensuring compatibility with spike-based encoding.
+The normalized skeleton input is then processed by a one-dimensional convolution (Conv1D), followed by batch normalization (BN) and a spiking neuron (SN) layer.
+This sequence forms the initial encoding stage of Signal-SGN, and the resulting feature is denoted as $X_{\text{i}}$.
+Figure 3. 3D skeleton visualization showing both positive and negative coordinate values.
+3.2. 1D-SGC Module
 In the 1D-SGC module, inspired by GCN, we integrate topological information by normalizing adjacency-based weights to aggregate attributes from neighboring nodes, as shown in Figure 2(a).
 The skeleton can be represented as a graph $\mathcal{G}(V,E)$, where the joints form a set of vertices $V$ and the bones represent the edges $E$.
-This structure can be expressed as an adjacency matrix $A\in\mathbb{R}^{V\times V}$, where $A_{i,j}=1$ if joints $i$ and $j$ are physically connected, and $A_{i,j}=0$ otherwise.
+This structure can be expressed as an adjacency matrix $A\in\mathbb{R}^{V\times V}$, where $A_{i,j}=1$ if joints $i$ and $j$ are physically connected, otherwise $A_{i,j}=0$.
 To incorporate topological relationships into computations, the normalized adjacency matrix $S$ is computed as follows:
-$S=\Delta^{-\frac{1}{2}}\widetilde{A}\Delta^{-\frac{1}{2}},$ (2)
-where $\widetilde{A}=A+I$ is the adjacency matrix with added self-connections, and $\Delta$ is the degree matrix of $\widetilde{A}$.
+(3) $S=\Delta^{-\frac{1}{2}}\widetilde{A}\Delta^{-\frac{1}{2}},$
+where $\widetilde{A}=A+I$ is the adjacency matrix with added self-connections $I$, and $\Delta$ is the degree matrix of $\widetilde{A}$.
 This normalization ensures that the influence of each node is appropriately scaled based on its connectivity, facilitating the effective integration of topological information into the model.
-Subsequently, we apply one-dimensional convolution (Conv1d) and one-dimensional Batch Normalization(BN) along the $C^{\prime}$ dimension to extract features for each skeleton node within each spike time step, effectively aggregating features across the skeleton joints dimension.
-$G_{\text{i}}=\mathrm{BN}\left[\sum_{n=0}^{N}\operatorname{Conv1d}\left(X_{\text{i}}*S^{n}\right)\right],$ (3)
+Subsequently, we apply Conv1d and BN along channel dimension to extract features for each skeleton node within each spike time step, effectively aggregating features across the skeleton joints dimension.
+(4) $G_{\text{i}}=\mathrm{BN}\left[\sum_{n=0}^{N}\operatorname{Conv1d}\left(X_{\text{i}}*S^{n}\right)\right],$
 in which $S^{n}$ represents a set of matrices decomposed from $S$, $N$ refers to the number of different relational matrices in the decomposition, and * denotes matrix multiplication.
@@
 […]
-$G^{(0)}_{\text{o}}=SN\left(G_{\text{i}}\right)+\Phi\left(X_{\text{i}}\right).$ (4)
-Building upon the approach proposed in [13], we integrate the spike self-attention (SSA) mechanism after the graph convolution operation to enhance feature representation. The formula is as follows:
-$Q^{(0)}=\Phi_{Q}(G^{(0)}_{\text{o}}),K^{(0)}=\Phi_{K}(G^{(0)}_{\text{o}}),V^{(0)}=\Phi_{V}(G^{(0)}_{\text{o}}),$ (5)
-$G^{(0)}=G^{(0)}_{\text{o}}+\Phi\left(Q^{(0)}K^{(0)T}V^{(0)}*s\right),$ (6)
+(5) $G^{(0)}_{\text{o}}=SN\left(G_{\text{i}}\right)+\Phi\left(X_{\text{i}}\right).$
+Building upon the approach proposed in (Zhou et al., 2022), we integrate the spike self-attention (SSA) mechanism after the graph convolution operation to enhance feature representation.
+The formula is as follows:
+(6) $Q^{(0)}=\Phi_{Q}(G^{(0)}_{\text{o}}),K^{(0)}=\Phi_{K}(G^{(0)}_{\text{o}}),V^{(0)}=\Phi_{V}(G^{(0)}_{\text{o}}),$
+(7) $G^{(0)}=G^{(0)}_{\text{o}}+\Phi\left(Q^{(0)}K^{(0)T}V^{(0)}*s\right),$
 where $s$ is the scaling factor to control the large value of the matrix multiplication result.
+3.3. FSC Module
 Following the 1D-SGC module, the features are re-transformed by $SN$ into spiking-form feature—discrete, aperiodic signal sequences encoded as binary values (0s and 1s).
 To further extract meaningful frequency-domain feature, we propose the FSC module, as shown in Figure 2(b).
-This module first leverages a learnable window function applied along the joint dimension $V$ to mitigate spectral leakage, enhancing the precision of frequency-domain analysis and improving feature extraction robustness, defined as:
-$w_{V}(v)=\alpha-(1-\alpha)\cdot\cos\left(\frac{2\pi v}{V-1}\right),\quad 0\leq v\leq V-1.$ (7)
-where $v$ is the joint index and $\alpha$ is a learnable parameter controlling the window shape.
-The learnability of $\alpha$ allows the module to adapt the window function to specific characteristics.
-The windowed features are then processed as:
-$F^{(0)}_{\text{i}}=w_{V}(v)\cdot SN(G^{(0)})\in\mathbb{R}^{T\times D\times V}.$ (8)
-We apply FFT along the $V$ dimension to transform the features from the temporal domain to the frequency domain:
-$F^{(0)}_{\text{o}}=\sum_{v=0}^{V-1}F^{(0)}_{\text{i}}e^{-j2\pi\frac{mv}{V}}=F^{(0)}_{\text{R}}+jF^{(0)}_{\text{I}},$ (9)
-where $F^{(0)}_{\text{o}}\in\mathbb{C}^{T\times D\times M}$, and $m\in[0,M-1]$ represents the frequency index, determining the frequency components extracted along the $V$ dimension.
+First, this module leverages a learnable window function $w_{V}(v)\in\mathbb{R}^{V}$ along the joint dimension to mitigate spectral leakage and enhance frequency analysis.
+Unlike fixed windows, $w_{V}$ assigns adaptive weights to each joint, acting as a soft spatial gate before the frequency transform and applied element-wise to the input feature map $G^{(0)}$ as:
+(8) $F^{(0)}=G^{(0)}\cdot w_{V}(v),$
+where $F^{(0)}$ denotes the windowed features passed to the frequency encoder, and $\cdot$ denotes broadcasting over the joint dimension.
+We perform a Fast Fourier Transform (FFT) along the joint dimension $V$ to project the input feature $F^{(0)}_{\text{i}}$ into the frequency domain.
+The resulting complex-valued representation is separated into real and imaginary components, denoted as $F^{(0)}_{\text{R}}$ and $F^{(0)}_{\text{I}}$, which are subsequently processed by parallel branches.
+This transformation enables the model to extract interpretable spatial-frequency patterns across joints, facilitating enhanced action representation.
 To process the complex-valued FFT outputs, we propose a spiking complex convolution module that preserves and utilizes amplitude and phase information, enhancing the network’s ability to capture detailed frequency-domain features:
-$F^{(0)}_{\text{o},\text{R}}=\mathrm{BN}[\operatorname{Conv1d}(F^{(0)}_{\text{R}})],F^{(0)}_{\text{o},\text{I}}=\mathrm{BN}[\operatorname{Conv1d}(F^{(0)}_{\text{I}})],$ (10)
-$F^{(0)}_{\text{o},k}=\text{SN}\left(F^{(0)}_{\text{o},\text{R}}\pm_{k}F^{(0)}_{\text{o},\text{I}}\right),\quad k\in\{1,2\}.$ (11)
-where ±$\pm_{k}$ represents addition ($+$) when $k=1$, and subtraction (--) when $k=2$. The fused output is then computed as:
-$F^{(0)}=\Phi\left(F^{(0)}_{\text{o},1}\right)+\Phi\left(F^{(0)}_{\text{o},2}\right).$ (12)
+(9) $F^{(0)}_{\text{o},\text{R}}=\mathrm{BN}[\operatorname{Conv1d}(F^{(0)}_{\text{R}})],F^{(0)}_{\text{o},\text{I}}=\mathrm{BN}[\operatorname{Conv1d}(F^{(0)}_{\text{I}})],$
+(10) $F^{(0)}_{\text{o},u}=\text{SN}\left(F^{(0)}_{\text{o},\text{R}}\pm_{u}F^{(0)}_{\text{o},\text{I}}\right),\quad u\in\{1,2\}.$
+where ±$\pm_{k}$ represents addition ($+$) when $u=1$, and subtraction (--) when $u=2$. The fused output is then computed as:
+(11) $F^{(0)}=\Phi\left(F^{(0)}_{\text{o},1}\right)+\Phi\left(F^{(0)}_{\text{o},2}\right).$
 The output $F^{(0)}$ represents the extracted frequency-domain features, serving as an essential input for the subsequent layers of the network.
-A complete 1D-SGC-FSC process can then be expressed as:
-$X^{(l)}=F^{(l)}+G^{(l)}+\operatorname{res}\left(G^{(l-1)}\right),$ (13)
-where, $X^{(l)}$ denotes the output of the $l$-th layer, combining the frequency-domain features $F^{(l)}$, temporal-domain features $G^{(l)}$, and a residual connection from the previous layer $\operatorname{res}\left(G^{(l-1)}\right)$.
-III-B Multi-Scale Wavelet Transform Feature Fusion module
+Then a complete 1D-SGC-FSC process can then be expressed as:
+(12) $X^{(l)}=F^{(l)}+G^{(l)}+\operatorname{res}\left(G^{(l-1)}\right)$
+where $X^{(l)}$ denotes the output of the $l$-th layer, combining the frequency-domain features $F^{(l)}$, temporal-domain features $G^{(l)}$, and a residual connection from the previous layer $\operatorname{res}\left(G^{(l-1)}\right)$.
+3.4. Multi-Scale Wavelet Transform Feature Fusion module
 We formulate the MWTF module to capture the temporal dynamics and frequency-specific characteristics of spiking-form features, with its outputs fused with backbone outputs to enhance classification performance, as illustrated in Figure 2(c).
-We first construct filters $\Lambda_{0}$, $\Lambda_{1}$, $\Gamma_{0}$, and $\Gamma_{1}$ based on Legendre polynomials $P_{k}(x)$ (details are provided in Appendix V-A) to decompose spiking-form features.
+We first construct filters $\Lambda_{0}$, $\Lambda_{1}$, $\Gamma_{0}$, and $\Gamma_{1}$ based on Legendre polynomials $P_{k}(x)$ (details are provided in Appendix B) to decompose spiking-form features.
 The parameter $k$ specifies the number of frequency coefficients extracted, influencing the spectral resolution.
-We apply a linear (LN) transformation to the spiking features from the final layer $L$ of the backbone network to align the output tensor with the $k$ coefficients, ensuring compatibility with the filters.
-$X_{\text{down}}=\operatorname{LN}\left(SN\left(X^{(L)}\right)\right)\in\mathbb{R}^{T\times k\times V}.$ (14)
-To ensure compatibility with the wavelet transform, we verify whether the temporal dimension $T$ is a power of two. If not, $X_{\text{down}}$ is padded to the nearest power of two:
-$X_{\text{extra}}=X_{\text{down}}[0:(2^{\lceil\log_{2}(T)\rceil}-T),:,:],$ (15)
-$X_{\text{extended }}=\operatorname{concat}(X_{\text{down}},X_{\text{extra}},\text{ axis }=1),$ (16)
-The extended spike-form feature $X_{\text{extended }}$ is decomposed by the filter matrices $\Lambda_{0}$, $\Lambda_{1}$, $\Gamma_{0}$, and $\Gamma_{1}$ into detail coefficients $D_{j}$, capturing high-frequency detailed features, and scaling coefficients $S_{j}$, retaining low-frequency trend features.
-The Algorithm 2 in Appendix V-B2 details the complete process.
-To enhance the interaction between detail coefficients $D_{j}$ and scaling coefficients $S_{j}$ across decomposition levels $j$, we redefine the spiking cross-attention operation as:
-$D_{j}^{\text{out}}=\left(\Phi_{Q}(D_{j})\cdot\Phi_{K}(D_{j})^{T}\right)\Phi_{V}(S_{j})*s,$ (17)
-$S_{j}^{\text{out}}=\left(\Phi_{Q}(S_{j})\cdot\Phi_{K}(S_{j})^{T}\right)\Phi_{V}(D_{j})*s,$ (18)
-where the temporal resolution is halved at each level.
-At the final level $J$, the scaling coefficient $S_{J}$ bypasses cross-attention and is directly upsampled to match the temporal resolution $T/2$:
-$\hat{S}_{J}=\text{Upsample}(S_{J},T/2).$ (19)
-All previous levels’ outputs $\{D_{j}^{\text{out}},S_{j}^{\text{out}}\}_{j=0}^{J-1}$ are similarly upsampled, and the aggregated feature map is computed as:
-$X_{\text{MLO}}=\operatorname{GAP}\left(\Phi\left(\sum_{j=0}^{J-1}\hat{D}_{j}+\sum_{j=0}^{J}\hat{S}_{j}\right)\right).$ (20)
-The final feature map $X_{\text{MLO}}$ integrates dynamic spiking features and spectral characteristics, which are fused with the output $X^{(L)}$ from the last backbone layer undergoing $SN$ and GAP.
-Then the fused feature proceed to a fully connected (FC) layer for classification:
-$\hat{y}=\operatorname{FC}\left(\operatorname{GAP}\left(SN\left(X^{(L)}\right)\right)+\beta X_{\text{MLO}}\right),$ (21)
-where $\beta$ is a learnable parameter. Finally, the cross-entropy loss is employed to optimize the classification task.
-IV Experiments
-We evaluate the Signal-SGN model on three skeleton-based action recognition datasets: NTU-RGB+D [30], NTU-RGB+D 120 [31], and NW-UCLA [32].
-Detailed descriptions of the datasets, along with hyperparameter design, FLOPs, synaptic operations (SOPs), and theoretical energy consumption calculations, are provided in the Appendix V-C, V-D and V-E.
-IV-A Comparison with the State-of-the-Art
-TABLE I: Comparative results on NTU RGB+D and NTU-RGB+D 120. We evaluate our model in terms of classification accuracy (%). Xs and Xv represent cross-subject and cross-view splits.
-Model (ANN/SNN) Param. (M) FLOPs (G) SOPs (G) Power (mJ) Xs (60) Xv (60) Xs (120) Xv (120)
-Part-aware LSTM [33] - - - - 62.9 70.3 25.5 26.3
-Spatio-Temporal LSTM [34] - - - - 61.7 75.50 55.7 57.9
-ST-GCN [6] 3.10 3.48 - 16.01 81.5 88.3 70.7 73.2
-2S-AGCN [7] 3.48 37.32 - 182.87 88.5 95.1 82.5 84.3
-Shift-GCN [5] - 10.00 - 46.0 90.7 96.5 85.3 86.6
-MS-G3D [35] 3.19 48.88 - 239.51 91.5 96.2 86.9 88.4
-CTR-GCN [9] 1.46 7.88 - 36.25 92.4 96.8 88.9 90.6
-Spikformer [13] ICLR 2023 4.78 24.07 1.69 2.17 73.9 80.1 61.7 63.7
-Spike-driven Transformer [14] NIPS 2023 4.77 23.50 1.57 1.93 73.4 80.6 62.3 64.1
-Spike-driven Transformer V2 [15] ICLR 2024 11.47 38.28 2.59 2.91 77.4 83.6 64.3 65.9
-Spiking Wavelet Transformer [29] ECCV 2025 3.24 19.3 1.48 2.01 74.7 81.2 63.5 64.7
-MK-SGN [17] 2.17 7.84 0.68 0.614 78.5 85.6 67.8 69.5
-Signal-SGN (Joint) 1.74 1.62 0.314 0.372 80.5 87.7 69.2 72.1
-Signal-SGN (Bone+Joint) 1.74 3.24 0.628 0.744 82.5 89.2 71.3 74.2
-Signal-SGN (4 ensemble) 1.74 6.48 1.28 1.488 86.1 93.1 75.3 77.9
-The Param.
+We apply a linear transformation to the spiking features from the final layer $L$ of the backbone network to align the output tensor with the $k$ coefficients, ensuring compatibility with the filters.
+(13) $X_{\text{down}}=\operatorname{LN}\left(SN\left(X^{(L)}\right)\right)\in\mathbb{R}^{T\times k\times V}.$
+To ensure compatibility with the wavelet transform, we first check whether the temporal dimension $T$ is a power of two.
+If it is not, we extend the input spike-form feature $X_{\text{down}}$ by concatenating extra frames, denoted as $X_{\text{extra}}$, which are duplicated copies of the beginning frames.
+The total number of time steps is then increased to the nearest power of two.
+Formally, the number of extra frames added is calculated as $2^{\lceil\log_{2}T\rceil}-T$, and the final extended input becomes $\widetilde{X}=[X_{\text{down}},X_{\text{extra}}]$ along the temporal axis.
+The extended input $\widetilde{X}$ is then recursively decomposed by a set of wavelet filter matrices, namely $\Lambda_{0}$, $\Lambda_{1}$, $\Gamma_{0}$, and $\Gamma_{1}$, yielding two types of outputs at each level: detail coefficients $D_{j}$, which capture high-frequency, localized temporal features, and scaling coefficients $S_{j}$, which preserve the lower-frequency trend components, as shown in Figure 4.
+A complete iterative decomposition process is provided in Algorithm 2 in Appendix C.2.
+Figure 4. The process of DWT
+To enhance the interaction between detail coefficients $D_{j}$ and scaling coefficients $S_{j}$ across decomposition levels $j$, we first upsample them to the original temporal resolution $T$.
+Then, a spiking cross-attention mechanism is applied to capture the mutual dependencies between high-frequency and low-frequency components:
+(14) $\hat{D}_{j}=\left(\Phi_{Q}(D_{j})\cdot\Phi_{K}(D_{j})^{T}\right)\Phi_{V}(S_{j})\cdot s,$
+(15) $\hat{S}_{j}=\left(\Phi_{Q}(S_{j})\cdot\Phi_{K}(S_{j})^{T}\right)\Phi_{V}(D_{j})\cdot s,$
+where $s$ is a learnable scaling factor.
+The scaling coefficient from the final level $S_{J}$ is only upsampled (without interaction) and denoted as $\hat{S}_{J}$.
+All outputs $\{\hat{D}_{j},\hat{S}_{j}\}_{j=0}^{J-1}\cup\{\hat{S}_{J}\}$ are aggregated and passed through spiking activation and global average pooling to yield the multi-level output:
+(16) $\hat{X}=\operatorname{GAP}\left(\Phi\left(\sum_{j=0}^{J-1}\hat{D}_{j}+\sum_{j=0}^{J}\hat{S}_{j}\right)\right).$
+The resulting feature $\hat{X}$ is then fused with the SN-processed output of the final backbone layer $X^{(L)}$, and the fused representation is sent to a fully connected layer for classification:
+(17) $\hat{y}=\operatorname{FC}\left(\operatorname{GAP}\left(SN\left(X^{(L)}\right)\right)+\beta\hat{X}\right),$
+where $\beta$ is a learnable parameter balancing the contribution of temporal and spectral features. Cross-entropy loss is used for optimization during training.
+4. Experiments
+To demonstrate the effectiveness of Signal-SGN, we conduct skeleton-based action recognition experiments on three large-scale datasets.
+We compare our model against strong ANN and SNN baselines, and perform ablation studies to evaluate the contribution of each module.
+The following sections cover the experimental setup in detail, including dataset settings, hyperparameter configurations, comparisons with state-of-the-art methods, and ablation studies.
+4.1. Dataset
+Table 1.
+Comparison of Signal-SGN with prior ANN- and SNN-based models on NTU RGB+D, NTU RGB+D 120, and NW-UCLA datasets.
+We report classification accuracy (%), model size (Param.), computational cost (FLOPs and SOPs), and energy consumption (Power).
+Model (ANN/SNN) Param. FLOPs SOPs Power NTU RGB+D NTU RGB+D 120 NW-UCLA
+(M) (G) (G) (mJ) Xs(%) Xv(%) Xs(%) Xt(%) (%)
+Part-aware LSTM (Du Y., 2015) - - - - 62.9 70.3 25.5 26.3 -
+Spatio-Temporal LSTM (Liu et al., 2016) - - - - 61.7 75.5 55.7 57.9 -
+ST-GCN (Yan et al., 2018) 3.10 3.48 - 16.01 81.5 88.3 70.7 73.2 -
+2S-AGCN (Shi et al., 2019) 3.48 37.32 - 182.87 88.5 95.1 82.5 84.3 -
+Shift-GCN (Cheng et al., 2020b) - 10.00 - 46.00 90.7 96.5 85.3 86.6 94.6
+MS-G3D (Liu et al., 2020) 3.19 48.88 - 239.51 91.5 96.2 86.9 88.4 -
+CTR-GCN (Chen et al., 2021) 1.46 7.88 - 36.25 92.4 96.8 88.9 90.6 96.5
+Spikformer (Zhou et al., 2022) 4.78 24.07 1.69 2.17 73.9 80.1 61.7 63.7 85.4
+Spike-driven Transformer (Yao et al., 2024b) 4.77 23.50 1.57 1.93 73.4 80.6 62.3 64.1 83.4
+Spike-driven Transformer V2 (Yao et al., 2024a) 11.47 38.28 2.59 2.91 77.4 83.6 64.3 65.9 89.4
+Spiking Wavelet Transformer (Fang et al., 2025) 3.24 19.30 1.48 2.01 74.7 81.2 63.5 64.7 86.7
+MK-SGN (Zheng et al., 2024) 2.17 7.84 0.68 0.614 78.5 85.6 67.8 69.5 92.3
+Signal-SGN (Joint) 1.74 1.62 0.314 0.372 80.5 87.7 69.2 72.1 92.7
+Signal-SGN (Bone+Joint) 1.74 3.24 0.628 0.744 82.5 89.2 71.3 74.2 93.1
+Signal-SGN (4 ensemble) 1.74 6.48 1.28 1.488 86.1 93.1 75.3 77.9 95.9
+Param.
 represents the model size for the network itself.
-For ANN, the metrics such as FLOPs, power, and accuracy are reported and calculated based on the maximum ensemble of modalities as specified in the referenced paper.
-TABLE II: Comparative results on NW-UCLA. We evaluate our model in terms of classification accuracy (%).
-Model Param. FLOPs SOPs Power Acc
-(ANN/SNN) (M) (G) (G) (mJ) (%)
-Ensemble TS-LSTM [3] - - - - 89.2
-DC-GCN+ADG [18] 2.48 3.6 - 16.56 95.3
-CTR-GCN [9] 1.46 2.32 - 10.67 96.5
-Spikformer [13]ICLR 2023 4.78 8.55 0.513 0.673 85.4
-Spike-driven Transformer [14]NIPS 2023 4.77 8.22 0.501 0.643 83.4
-Spike-driven Transformer V2 [15]ICLR 2024 11.47 11.96 0.717 0.841 89.4
-Spiking Wavelet Transformer [29]NIPS 2023 3.24 7.36 0.438 0.573 86.7
-MK-SGN [17] 2.17 3.66 0.165 0.207 92.3
-Signal-SGN (Joint) 1.74 0.67 0.161 0.147 92.7
-Signal-SGN (Bone+Joint) 1.74 1.24 0.322 0.294 93.1
-Signal-SGN (4 ensemble) 1.74 2.48 0.644 0.588 95.9
-Several SOTA methods have adopted multi-stream fusion strategies [7, 5, 9].
-For a fair comparison, we implement the same four-stream fusion approach for our Signal-SGN, incorporating ”joint”, ”bone,” ”joint motion”, and ”bone motion.” We compare Signal-SGN with state-of-the-art methods across three datasets.
-To compare with SOTA SNN models[13, 14, 15, 29], we perform unimodal reproductions using the same hierarchical architecture as MK-SGN[17], demonstrating the superiority of our method.
-The parameters of our Signal-SGN model, including FLOPs, SOPs, energy consumption, and accuracy for both single-modal and multi-modal scenarios, are presented in Tables I and II.
-In the NTU-RGB+D dataset, we resize the skeleton sequences to 16 frames, our model outperforms SOTA SNN models in accuracy while significantly reducing FLOPs, SOPs, and theoretical energy consumption when using single-modality joint data.
-Furthermore, compared to classical ANN models, our model achieves a higher accuracy than ST-GCN when integrating four modalities, which is quite rare for SNN.
-Our model also requires fewer FLOPs than the lightweight CTR-GCN and Shift-GCN models.
-In terms of theoretical energy consumption, our model reduces the energy usage to less than 5% compared to GCN-based methods, which is highly significant in practical applications.
-In the NW-UCLA dataset, we also resize the skeleton sequences to 16 frames.
-Our model achieves an accuracy of 95.9%, which is only 0.6% lower than the classic CTR-GCN model while reducing theoretical energy consumption to 5.51% of CTR-GCN’s.
-This further demonstrates the effectiveness and energy-saving capabilities of our model.
-Additionally, removing the extra spike time step dimension reduces FLOPs and improves memory efficiency, making the model more trainable on standard devices compared to the previous SGN model.
-IV-B Ablation Studies
-To evaluate the contributions of the primary components in Signal-SGN, we conducted ablation studies on the NTU-RGB+D (cross-subject) dataset using joint data, with all skeleton sequences resized to 16 frames.
-Table III summarizes classification performance for different main module configurations.
-Appendix V-F presents a detailed analysis.
-TABLE III: Comparison of experimental results under different module settings
-Method 1D-SGC FSC MWTF Param. FLOPs/SOPs Acc
-(M) (G) (%)
-Ours - - 1.03 1.12/0.252 70.5
-- 1.65 1.60/0.312 78.3
-1.74 1.62/0.314 80.5
-When only the 1D-SGC module is used, the model achieves an accuracy of 70.5%, demonstrating its ability to extract spiking-form graph features while maintaining low parameter count (1.03M) and computational costs (1.12G FLOPs/0.252G SOPs).
-Introducing FSC module significantly improves accuracy to 78.3%, highlighting the role of frequency-domain feature extraction in enhancing representations.
-The addition results in a moderate increase in parameter count to 1.65M and computational costs to 1.60G FLOPs/0.312G SOPs.
-Finally, incorporating the MWTF module further boosts accuracy to 80.5%.
-The improvement comes with a marginal parameter increase (1.74M) and computational costs (1.62G FLOPs/0.314G SOPs), demonstrating the MWTF module’s effectiveness in capturing multi-scale frequency features and fusing temporal and spectral information.
-These results validate the complementary roles of the 1D-SGC, FSC, and MWTF modules.
-The 1D-SGC module focuses on temporal-domain graph features, while the FSC module captures frequency-domain information.
-The MWTF module further enhances the model by integrating multi-scale frequency features, establishing Signal-SGN as an efficient framework for skeleton-based action recognition.
-V Conclusion
+For ANN methods, FLOPs, power, and accuracy are reported under standard protocols.
+For SNN methods, SOPs and energy estimates are computed based on spike-driven inference.
+NTU RGB+D (Shahroudy et al., 2016) is a large-scale benchmark dataset for 3D human action recognition, consisting of 56,880 skeleton-based video samples from 60 action classes.
+Each sample contains 3D coordinates of 25 human body joints captured by the Kinect v2 sensor.
+We follow the standard evaluation protocols, including cross-subject (Xs)—where training and testing subjects do not overlap—and cross-view (Xv)—where training and testing samples are captured from different camera angles.
+NTU-RGB+D 120 (Liu et al., 2019) extends NTU 60 to 120 action categories and 114,480 video samples, with more subjects and greater view diversity.
+Evaluation is performed under cross-subject (Xs) and cross-setup (Xt) protocols, where the latter refers to training and testing samples being collected under different environmental and camera configurations, posing a greater challenge for model generalization.
+NW-UCLA (Wang et al., 2014) contains 10 action classes and 1,494 samples captured from three different viewpoints.
+Each sample includes 3D skeletons with 20 joints.
+The dataset is designed to test model robustness under cross-view scenarios.
+4.2. Model Implementation and Configuration
+Our model is implemented using PyTorch [^1] https://github.com/rwightman/pytorch-image-models and SpikingJelly [^2] https://github.com/fangwei123456/spikingjelly, and all experiments are conducted on four NVIDIA V100 GPUs.
+The backbone network consists of 4 layers with input and output channel sizes of 3 and 64, 64 and 64, 64 and 128, 128 and 256.
+We use SGD optimizer with a learning rate of 0.1, decay rate of 0.1 at step 110, batch size of 64, training for 150 epochs with a dropout rate of 0.3, weight decay of 0.0001, and auxiliary loss weight $\omega$ set to 0.03.
+4.3. Comparison with the State-of-the-Art
+Several SOTA methods adopt multi-stream fusion strategies for improved performance (Shi et al., 2019; Cheng et al., 2020b; Chen et al., 2021).
+To ensure a fair comparison, we employ the same four-stream fusion protocol in Signal-SGN, incorporating joint, bone, joint motion, and bone motion modalities.
+We evaluate Signal-SGN against both ANN and SNN baselines on NTU RGB+D, NTU RGB+D 120, and NW-UCLA.
+Classification accuracy, model size, FLOPs, synaptic operations(SOPs), and estimated energy consumption are summarized in Table 1, with calculation details provided in Appendix A.
+On the NTU RGB+D and NTU RGB+D 120 datasets, all sequences are downsampled to 16 frames.
+Compared to prior SNNs, Signal-SGN (Joint) achieves 80.5% accuracy on NTU RGB+D (Xs) while requiring only 0.372 mJ of energy—an 82.9% reduction compared to Spikformer’s 2.17 mJ, and a 98.9% reduction compared to the ANN-based MS-G3D (239.51 mJ).
+Notably, our four-stream fused model further boosts accuracy to 86.1% (Xs) and 93.1% (Xv) on NTU RGB+D and 75.3% (Xs) / 77.9% (Xt) on NTU RGB+D 120, while still keeping energy usage at 1.488 mJ, which is over 96% lower than CTR-GCN (36.25 mJ) and 99.4% lower than MS-G3D.
+This demonstrates Signal-SGN’s superior trade-off between performance and efficiency.
+On the NW-UCLA, Signal-SGN achieves a competitive 95.9% accuracy, only 0.6% below CTR-GCN, while using less than 5% of its theoretical energy—further validating the model’s generalizability and low-cost design.
+4.4. Ablation Studies
+To evaluate the contributions of the primary components in Signal-SGN, we conducted ablation studies on the NTU-RGB+D (Xs) dataset based on joint data, with all sequences resized to 16 frames.
+Table 2. Module-wise ablation study on NTU-RGB+D (Xs). We report model size, computational cost, and top-1 accuracy. Spikformer is included as a baseline for external comparison.
+3D C.N. 1D-SGC FSC MWTF Param. (M) FLOPs / SOPs (G) Acc (%)
+Spikformer (baseline)(Zhou et al., 2022) 4.78 24.07 / 1.69 73.9
+- - - 1.03 1.12 / 0.252 69.7
+- - 1.65 1.60 / 0.312 77.0+7.3
+- 1.74 1.62 / 0.314 79.3 +2.3
+- - 1.03 1.12 / 0.252 70.5
+- 1.65 1.60 / 0.312 78.3+7.8
+1.74 1.62 / 0.314 80.5+2.2
+4.4.1. Module-wise Ablation Analysis
+We first module-wise ablation analysis to evaluate the effectiveness of the proposed components in Signal-SGN.
+As summarized in Table 2, the base model with only the 1D-SGC module achieves 69.7% accuracy.
+Adding the FSC significantly boosts performance to 77.0%, indicating the advantage of incorporating frequency-domain information.
+Further introducing the MWTF module improves accuracy to 79.3%, demonstrating its ability to capture fine-grained temporal-frequency dynamics.
+A slight improvement is also observed when applying 3D coordinate normalization across different configurations, suggesting it can further stabilize input distributions and facilitate downstream spiking representation learning.
+Overall, these results confirm the progressive effectiveness of the proposed modules, with MWTF contributing most significantly on top of FSC.
+4.4.2. Ablation study on the FSC module
+Table 3. Ablation study on the FSC module.
+Configuration $u$ Interaction Window Acc (%)
+Without FSC 70.5
+Real branch only - Learnable 72.3+1.8
+Imaginary branch only - Learnable 72.1-0.2
+Real + Imag (no fusion) - Learnable 73.6+1.5
+Only $u=1$ branch $u=1$ Learnable 75.4+1.8
+Only $u=2$ branch $u=2$ Learnable 74.9-0.5
+Full FSC $u=1,2$ Learnable 78.3+3.4
+Full FSC $u=1,2$ Hamming 78.0-0.3
+Full FSC $u=1,2$ None 77.8-0.2
+To validate the effectiveness of the proposed FSC module in enhancing spiking feature representation, we conduct comprehensive ablation studies, as shown in Table 3.
+Introducing only the real or imaginary branch improves performance from 70.5% to 72.3% and 72.1%, respectively, suggesting that both components contribute meaningfully.
+Combining them without interaction ($u$) further increases accuracy to 73.6% (+1.5%), confirming their complementarity.
+Enabling single interaction branches—either $u=1$ or $u=2$—leads to 75.4% and 74.9%, respectively, while using both ($u=\{1,2\}$) achieves the best performance at 78.3%, yielding a +3.4% gain over the best single-branch configuration and +7.8% over the baseline.
+Replacing the learnable window with a fixed Hamming window slightly reduces accuracy to 74.9%, and removing the window entirely yields 77.8%, showing the window’s limited but non-negligible contribution.
+These results demonstrate that the core benefit of FSC arises from complex-valued modeling and frequency-domain interaction mechanisms.
+To further understand the role of the learnable window, we conduct a visualization analysis across FSC layers, as detailed in Appendix D.2.
+4.4.3. Qualitative Visualization of FSC module
+To better understand the spiking neural behavior of our model, we visualize the joint-time ($V\times T$) spiking heatmaps from a representative “drinking water” sample, as shown in Figure 5.
+For each of the four FSC layers, we display activations from the real, imaginary, and fused branches, where grayscale intensity encodes the summed spike responses over the channel dimension.
+Figure 5. Spike response visualizations of FSC for the “drinking water” action sample.
+Figure 6.
+Visualization of upsampled detail coefficients $\hat{D}_{j}$ and scaling coefficients $\hat{S}_{j}$ across levels $j=0,1,2$, along with the final approximation coefficient $S_{3}$ in MWTF.
+Each heatmap shows the spike response intensity over joints and time steps, revealing the multi-scale temporal characteristics captured by the spiking cross-attention mechanism.
+The fused features consistently exhibit smoother and more structured spatiotemporal patterns, indicating that combining real and imaginary branches enhances feature stability and expressiveness.
+In contrast, the real branch shows sparser, localized firing, while the imaginary branch captures more dispersed activity.
+Importantly, in the final FSC layer, joints associated with the right arm—specifically the right elbow (10), wrist (11), and hand (12)—exhibit prominent activations, aligning with the expected motion of lifting a cup.
+Symmetric activity is also observed in joints 6–8 (left arm), reflecting bilateral arm coordination.
+These observations validate that FSC learns semantically meaningful joint dynamics, and that the fused representation captures discriminative patterns critical for accurate action recognition.
+4.4.4. Ablation study on the MWTF module.
+To further analyze the effect of the MWTF module, we conduct an ablation study on its two key hyperparameters: $k$ (number of frequency coefficients) and $J$ (number of decomposition levels).
+The results, presented in Table 4, evaluate the trade-off between model complexity and classification accuracy.
+Table 4. Comparison of experimental results under different MWTF settings ($k$ and $J$)
+MWTF $k$ $J$ Param. (M) FLOPs/SOPs (G) Acc (%)
+- - - 1.65 1.60/0.312 78.3
+4 1 1.68 1.61/0.313 78.9+0.6
+8 1 1.73 1.62/0.314 79.8+0.9
+8 2 1.74 1.62/0.315 80.3+0.5
+8 3 1.74 1.62/0.317 80.5+0.2
+16 3 1.83 1.64/0.320 80.5-
+Starting from the baseline model without MWTF ($78.3\%$), incorporating low-order configurations ($k{=}4,J{=}1$) leads to a modest gain ($+0.6\%$), with the best performance ($80.5\%$) achieved at $k{=}8,J{=}3$.
+This validates the importance of multi-level frequency decomposition.
+However, further increasing $k$ to 16 brings no improvement, indicating saturation.
+Table 5. Ablation study on spiking cross-attention and inclusion of $S_{J}$ in MWTF ($k{=}8,J{=}3$).
+Setting $\hat{D}_{j}$ Branch $\hat{S}_{j}$ Branch $\hat{S}_{J}$ Acc. (%)
+None (Baseline) 78.3
+Only $\hat{D}_{j}$ 78.9+0.6
+Only $\hat{S}_{j}$ 78.7-0.2
+Only $S_{J}$ 78.4-0.3
+$\hat{D}_{j}$ + $\hat{S}_{j}$ 79.8+0.4
+$\hat{S}_{j}$ + $S_{J}$ 79.2-0.6
+$\hat{D}_{j}$ + $S_{J}$ 79.4+0.2
+Full (Ours) 80.5+1.3
+To further dissect MWTF’s contributions, we ablate the two spiking cross-attention branches ($\hat{D}_{j}$, $\hat{S}_{j}$) and the use of the final approximation coefficient $\hat{S}_{J}$.
+Results in Table 5 show that each component brings incremental benefits, with $\hat{D}_{j}$ Branch alone improving performance to $78.9\%$, and the full configuration reaching $80.5\%$.
+To complement the analysis, we also investigate the evolution of the learnable fusion weight $\beta$ training.
+The results, including its stabilization behavior, are provided in Appendix D.1.
+4.4.5. Interpretability Analysis of MWTF via CAM Visualization
+To further validate the interpretability of the MWTF module, we visualize the joint-level Class Activation Maps (CAMs) derived from its output features for three representative actions: Drink Water, Stand Up,Reach into Pocket,Phone Call,Tear up Paper and Throw(Figure 7).
+These CAMs highlight the spatial regions most responsive in the model’s representation.
+In the Drink Water example, high activation appears near the right wrist and elbow, matching the expected movement pattern.
+In Stand Up, the model emphasizes lower body joints, while Reach into Pocket shows localized attention around the hip and wrist regions.
+These joint-specific activations are consistent with the underlying motion characteristics of each action.
+The results demonstrate that MWTF’s multi-scale frequency decomposition effectively captures action-relevant joint responses, confirming the strength of time-frequency analysis in spiking skeleton-based action recognition.
+4.4.6. Qualitative Visualization of Multi-Scale Coefficients
+To assess the impact of multilevel decomposition and spiking cross-attention in MWTF, we visualize the upsampled detail $\hat{D}_{j}$ and scaling $\hat{S}_{j}$ coefficients ($j=0,1,2$), along with the final approximation coefficient $S_{3}$, for a representative drinking water action (Figure 6).
+The heatmaps reveal that $\hat{D}_{j}$ emphasizes localized, high-frequency joint activations, while $\hat{S}_{j}$ captures smoother, global trends.
+Particularly, $S_{3}$ exhibits temporally stable responses, reflecting semantic patterns preserved through hierarchical decomposition.
+The complementarity between $\hat{D}_{2}$ and $\hat{S}_{2}$ further demonstrates the effectiveness of spiking cross-attention in disentangling multi-scale temporal dynamics for robust recognition.
+A more comprehensive spectral analysis based on joint-wise frequency decomposition is provided in Appendix D.3.
+(a) Drink Water
+(b) Stand Up
+(c) Reach into Pocket
+(d) Phone Call
+(e) Tear Up Paper
+(f) Throw
+Figure 7. Class Activation Maps (CAMs) based on MWTF outputs for three representative actions. Warmer colors indicate higher joint-level importance.
+5. Conclusion
 We propose a novel Signal-SGN, a SNN-based framework designed to learn dynamic features in both the temporal and frequency domains for skeleton-based action recognition.
@@
 The final model surpasses SOTA SNNs in various aspects and achieves accuracy comparable to traditional ANN models while maintaining extremely low energy consumption, demonstrating its effectiveness and efficiency in skeleton-based action recognition.
-Appendix
-In the appendix, we first provide the supplemental detail description of Legendre Polynomial-Based Filters and pseudocode of the proposed methods, followed by a detailed introduction to the dataset.
-Then, we describe the hyperparameter settings and training details.
-V-A Legendre Polynomial-Based Filters
+Appendix A Power Consumption
+We assume that multiply-and-accumulate (MAC) and accumulate (AC) operations are implemented on the 45 nm technology node with $E_{\text{MAC}}$ = 4.6 pJ and $E_{\text{AC}}$ = 0.9 pJ (Horowitz, 2014).
+The $\mathrm{FLOPs}$ for Signal-SGN and all reproduced models are calculated using FVCore [^3] https://github.com/facebookresearch/fvcore, providing accurate estimates for each model component.
+These FLOPs are then combined with the firing rate and spike time steps to compute SOPs as:
+(18) $\mathrm{SOPs}(l)=fr\times T\times\mathrm{FLOPs}(l),$
+where $fr$ is firing rate, $T$ is the spike time steps, $\operatorname{FLOPs}(l)$ refers to the floating-point operations of $l$, defined as the number of MAC operations performed in $l$ block or layer, and $\operatorname{SOPs}$ are the number of spike-based AC operations.
+The model’s theoretical energy consumption is described as, according to (Tavanaei et al., 2019; Kundu et al., 2021; Zheng et al., 2021; Panda et al., 2020; Yao et al., 2021; Zhou et al., 2022):
+(19) $\begin{gathered}E_{\text{Signal-SGN }}=E_{\text{EAC}}\times\mathrm{FL}_{\mathrm{SNN}\text{ Conv }}^{i}+\\ E_{\text{AC}}\times\left[\left(\sum_{n={\hat{n}}}^{N}\mathrm{SOP}_{\mathrm{SNN}\text{ Conv }}^{n}+\sum_{m=1}^{M}\mathrm{SOP}_{\mathrm{SNN}\text{ FC }}^{m}\right.\right.\\ \left.\left.+\sum_{l=1}^{L}\mathrm{SOP}_{\mathrm{SNN}\text{ SSA }}^{l}+\sum_{f=1}^{F}\mathrm{SOP}_{\mathrm{SNN}\text{ FFT }}^{f}\right)+\mathrm{SOP}_{\mathrm{SNN}\text{ MWTF }}\right]\end{gathered}$
+where $\mathrm{FL}_{\mathrm{SNN}\text{ Conv }}$ is the first layer to encode skeleton into spike-form.
+Then the SOPs of $N$ SNN Conv layers, $M$ SNN Fully Connected Layer (FC), $L$ SSA layer , $F$ FFT layer and MWTF are added together are added together and multiplied by $E_{\text{AC}}$.
+The theoretical energy consumption for ANN and SNN models is calculated as follows:
+For ANNs, the energy consumption of a block $b$ is given by:
+(20) $\text{Power}(b)=E_{\text{EAC}}\times\mathrm{FLOPs}(b).$
+For SNNs, the energy consumption of a block $b$ is given by:
+(21) $\text{Power}(b)=E_{\text{AC}}\times\mathrm{SOPs}(b).$
+Appendix B Legendre Polynomial-Based Filters
 Legendre polynomials $P_{k}(x)$, defined over the interval $[-1,1]$, provide a robust orthogonal basis for constructing frequency-domain filters aimed at decomposing spike-like features in temporal signals.
@@
 The polynomials are generated recursively using the relation:
-$P_{0}(x)=1,\quad P_{1}(x)=x,$ (22)
+(22) $P_{0}(x)=1,\quad P_{1}(x)=x,$
 $(k+1)P_{k+1}(x)$ $=(2k+1)xP_{k}(x)-kP_{k-1}(x).$
@@
 This property is particularly advantageous for designing a set of filters that can isolate specific frequency components with minimal interference.
-To apply these polynomials to time-domain signals sampled over $[0,T]$, we first normalize the time index:
-$u=\frac{t}{T},\quad u\in[0,1].$ (24)
+To apply these polynomials to temporal-domain signals sampled over $[0,T]$, we first normalize the time index:
+(24) $u=\frac{t}{T},\quad u\in[0,1].$
 We then map $[0,1]$ to $[-1,1]$:
-$x=2u-1.$ (25)
+(25) $x=2u-1.$
 This transformation aligns the original time domain with the canonical domain of the Legendre polynomials, ensuring straightforward evaluation of $P_{k}(x)$ at the corresponding sampling points.
@@
 Specifically, the low-pass filters are defined as:
-$\Lambda_{0}[k,t]=\sqrt{2k+1}\,P_{k}(2u-1),$ (26)
-$\Lambda_{1}[k,t]=\sqrt{2k+1}\,P_{k}(2(u+0.5)-1)=\sqrt{2k+1}\,P_{k}(2u).$ (27)
+(26) $\Lambda_{0}[k,t]=\sqrt{2k+1}\,P_{k}(2u-1),$
+(27) $\Lambda_{1}[k,t]=\sqrt{2k+1}\,P_{k}(2(u+0.5)-1)=\sqrt{2k+1}\,P_{k}(2u).$
 Here, $\Lambda_{0}[k,t]$ and $\Lambda_{1}[k,t]$ share the same polynomial family but are evaluated at $12u-$ 1 and $2u$, respectively, introducing a phase-like shift that distinguishes the two filters in their temporal and spectral responses.
@@
 Nonetheless, a representative scheme, consistent with the normalization above, is given by:
-$\Gamma_{0}[k,t]=\sqrt{2}\,\Lambda_{0}[k,t],$ (28)
-$\Gamma_{1}[k,t]=\sqrt{2}\,\Lambda_{1}[k,t].$ (29)
+(28) $\Gamma_{0}[k,t]=\sqrt{2}\,\Lambda_{0}[k,t],$
+(29) $\Gamma_{1}[k,t]=\sqrt{2}\,\Lambda_{1}[k,t].$
 Taken together, these four filters $(\Lambda_{0},\Lambda_{1},\Gamma_{0},\Gamma_{1})$ can decompose the signal into distinct frequency bands, enabling a nuanced analysis of spike-like features.
 By carefully mapping temporal samples into the Legendre polynomial domain, employing a normalized and shifted set of low-pass filters, and incorporating a high-pass counterpart, this approach provides a stable, orthogonal, and frequency-selective analytical framework.
-This framework facilitates effective characterization, improved interpretability, and subsequent analyses of time-varying signals in a multi-resolution context.
-V-B Algorithm
+This framework facilitates effective characterization, improved interpretability, and subsequent analyses of temporal-varying signals in a multi-resolution context.
+Appendix C Algorithm
 This section presents the essential pseudocode and detail description underpinning the computational framework discussed in this work.
-V-B1 Spiking Neuron Dynamics Simulation
+C.1. Spiking Neuron Dynamics Simulation
 The Spiking Neuron Dynamics describes the core mechanisms governing the temporal dynamics of spiking neurons within the network, outlined in Algorithm 1.
@@
 Specifically, the potential evolves according to the equation:
-$\tau\frac{dV(t)}{dt}=-\left(V(t)-V_{\text{rest}}\right)+R\cdot G_{i}(t),$ (30)
+(30) $\tau\frac{dV(t)}{dt}=-\left(V(t)-V_{\text{rest}}\right)+R\cdot G_{i}(t),$
 where $G_{i}(t)=G[t,:,:]$ is the input tensor for time step $t$.
@@
 Subsequent layers of spiking neurons follow similar dynamics, allowing the network to propagate spike-based representations through deeper processing stages.
-V-B2 Iterative Decomposition
+C.2. Iterative Decomposition
 The Iterative Decomposition algorithm describes the process of decomposing spike-form features into their frequency domain representations.
@@
 Algorithm 2 Iterative Decomposition
-1:  $X_{0}\leftarrow X_{\text{extended}}$
+1:  $\widetilde{X}_{0}\leftarrow\widetilde{X}$
 2:  for $j=0$ to $J-1$ do
-3:     $X_{\text{even}}^{(j)}\leftarrow X_{j}[::2,:,:]$
-4:     $X_{\text{odd}}^{(j)}\leftarrow X_{j}[1::2,:,:]$
+3:     $X_{\text{even}}^{(j)}\leftarrow\widetilde{X}_{j}[::2,:,:]$
+4:     $X_{\text{odd}}^{(j)}\leftarrow\widetilde{X}_{j}[1::2,:,:]$
 5:     $D_{j}\leftarrow\Gamma_{0}\cdot X_{\text{even}}^{(j)}+\Gamma_{1}\cdot X_{\text{odd}}^{(j)}$
 6:     $S_{j}\leftarrow\Lambda_{0}\cdot X_{\text{even}}^{(j)}+\Lambda_{1}\cdot X_{\text{odd}}^{(j)}$
-7:     $X_{j+1}\leftarrow S_{j}$
+7:     $\widetilde{X}_{j+1}\leftarrow S_{j}$
 8:  end for
-9:  $S_{J}\leftarrow X_{J}$
+9:  $S_{J}\leftarrow\widetilde{X}_{J}$
 10:  return $\{D_{j}\}_{j=0}^{J-1}$ and $\{S_{j}\}_{j=0}^{J}$
-The iterative decomposition process begins with the extended spike-form feature tensor $X_{\text{extended}}$, which serves as the initial input.
+The iterative decomposition process begins with the extended spike-form feature tensor $\widetilde{X}$, which serves as the initial input.
 The tensor is initialized as $X_{0}$, providing the starting representation for decomposition across $J$ levels.
@@
 These components are independently processed to compute the detail coefficients $D_{j}$ and scaling coefficients $S_{j}$ as follows:
-$D_{j}=\Gamma_{0}\cdot X_{\text{even}}^{(j)}+\Gamma_{1}\cdot X_{\text{odd}}^{(j)},$ (31)
-$S_{j}=\Lambda_{0}\cdot X_{\text{even}}^{(j)}+\Lambda_{1}\cdot X_{\text{odd}}^{(j)}.$ (32)
+(31) $D_{j}=\Gamma_{0}\cdot X_{\text{even}}^{(j)}+\Gamma_{1}\cdot X_{\text{odd}}^{(j)},$
+(32) $S_{j}=\Lambda_{0}\cdot X_{\text{even}}^{(j)}+\Lambda_{1}\cdot X_{\text{odd}}^{(j)}.$
 where, $\Gamma_{0},\Gamma_{1}$ are high-pass filter matrices used to extract the detail coefficients $D_{j}$, which capture high-frequency temporal variations.
@@
 Upon completion, the algorithm outputs a set of detail coefficients $\{D_{j}\}_{j=0}^{J-1}$ and the final scaling coefficients $S_{J}$, collectively encapsulating the hierarchical temporal structure of the input.
-The complete iterative decomposition process is outlined in Algorithm 2, providing a systematic and computationally efficient method for processing spike-form freature.
-V-C Datasets
-V-C1 NTU-RGB+D
-NTU-RGB+D[30] is a comprehensive dataset widely used in the field of skeleton-based action recognition.
-It contains 60 action classes and 56,000 video clips, making it a robust benchmark for evaluating action recognition models.
-The dataset provides detailed 3D skeleton data, RGB frames, depth maps, and infrared sequences, covering a diverse range of daily activities such as eating, walking, and interacting with objects.
-The skeleton data consists of 25 joints per subject, captured using the Microsoft Kinect v2 sensor, which allows for precise modeling of human movement.
-This dataset is particularly valuable for researchers because it includes a variety of subjects performing actions from multiple viewpoints, offering a rich source of information to develop and test models capable of understanding complex human behaviors.
-V-C2 NTU-RGB+D 120
-NTU-RGB+D 120[31] is an extended version of the original NTU-RGB+D dataset, expanding the number of action classes from 60 to 120 and increasing the total number of video clips to over 114,000.
-This extension includes additional action categories that encompass more complex and subtle human interactions, such as ”taking a selfie” and ”writing on a board.” With 106 distinct subjects and more diverse camera angles, NTU-RGB+D 120 offers a more challenging and comprehensive benchmark for evaluating the scalability and generalization capabilities of action recognition models.
-The dataset retains the detailed skeleton tracking of its predecessor, providing 25 joints per person and ensuring that the data remains highly informative for developing models that require an in-depth understanding of spatial-temporal dynamics in human actions.
-As one of the largest skeleton-based action recognition datasets available, NTU-RGB+D 120 is instrumental in pushing the boundaries of research in this field.
-V-C3 NW-UCLA
-The NW-UCLA dataset[32] is a smaller but uniquely challenging dataset consisting of 10 action classes and 1,494 video clips.
-It is specifically designed to test the robustness of action recognition models against viewpoint variations.
-The dataset features actions like jumping, waving, and running, captured from multiple viewpoints using three Kinect cameras placed at different angles.
-Each action is performed by 10 subjects, offering diverse skeleton data with 20 joints per person.
-This setup mimics real-world scenarios where actions are observed from different perspectives, making it a critical dataset for evaluating how well models can maintain performance when faced with varying observational conditions.
-The NW-UCLA dataset is ideal for researchers aiming to enhance the adaptability and resilience of their models to changes in camera viewpoints, which is essential for applications in surveillance, human-computer interaction, and robotics.
-V-D Model Implementation and Configuration
-The models used for conducting experiments are implemented using PyTorch [^1] https://github.com/rwightman/pytorch-image-models, a classical deep learning framework known for its flexibility and efficiency, and SpikingJelly [^2] https://github.com/fangwei123456/spikingjelly, a specialized library for spiking neural network research.
-Our implementation leverages the powerful features of these tools to effectively model and simulate the complex dynamics of spiking neural networks.
-The training process is conducted on four NVIDIA V100 GPUs, which provide the necessary computational power and memory to handle the large datasets and intricate computations involved in our experiments.
-This setup ensures that our models are trained efficiently, allowing us to explore the full potential of our approach in a scalable and robust manner.
-TABLE IV: The parameter Settings of the Backbone
-Level of Layers Channel Size
-Input Output
-1 3 64
-2 64 64
-3 64 128
-4 128 256
-In this section, we also give the specific hyperparameters of the model and training settings in all experiments.
-Table IV.
-presents the parameter settings of the backbone network, detailing the input and output channel sizes across different layers to illustrate the model’s architectural configuration.
-The various hyperparameter settings used in our experiments are also shown in the TableV.
-TABLE V: Hyper-parameter training settings of Signal-SGN.
-Parameter NTU-RGB+D NTU-RGB+D NW-UCLA
-60 120
-Learning Rate 0.1/0.1 0.1/0.1 0.1/0.1
-Decay rate 0.1/0.1 0.1/0.1 0.1/0.1
-Batch Size 64 64 128
-Training Epochs 90 90 90
-Step 50 50 50
-Dropout Rate 0.3 0.3 0.3
-Weight Decay 1e-4 1e-4 1e-4
-$\omega$ 0.03 0.03 0.03
-Optimizer SGD SGD SGD
-V-E Power Consumption
-We assume that multiply-and-accumulate (MAC) and accumulate (AC) operations are implemented on the 45 nm technology node with $E_{MAC}$ = 4.6 pJ and $E_{AC}$ = 0.9 pJ.
-The $\mathrm{FLOPs}$ for Signal-SGN and all reproduced models are calculated using FVCore [^3] https://github.com/facebookresearch/fvcore, providing accurate estimates for each model component.
-These FLOPs are then combined with the firing rate and spike time steps to compute SOPs as:
-$\mathrm{SOPs}(l)=fr\times T\times\mathrm{FLOPs}(l),$ (33)
-where $fr$ is firing rate, $T$ is the spike time steps, $\operatorname{FLOPs}(l)$ refers to the floating-point operations of $l$, defined as the number of MAC operations performed in $l$ block or layer, and $\operatorname{SOPs}$ are the number of spike-based AC operations.
-The model’s theoretical energy consumption is described as:
-$\begin{gathered}E_{\text{Signal-SGN }}=E_{EAC}\times\mathrm{FL}_{\mathrm{SNN}\text{ Conv }}^{1}+\\ E_{AC}\times\left[\left(\sum_{n=2}^{N}\mathrm{SOP}_{\mathrm{SNN}\text{ Conv }}^{n}+\sum_{m=1}^{M}\mathrm{SOP}_{\mathrm{SNN}\text{ FC }}^{m}\right.\right.\\ \left.\left.+\sum_{l=1}^{L}\mathrm{SOP}_{\mathrm{SNN}\text{ SSA }}^{l}+\sum_{f=1}^{F}\mathrm{SOP}_{\mathrm{SNN}\text{ FFT }}^{f}\right)+\mathrm{SOP}_{\mathrm{SNN}\text{ MWTF }}\right]\end{gathered}$ (34)
-where $\mathrm{FL}_{\mathrm{SNN}\text{ Conv }}^{1}$ is the first layer to encode skeleton into spike-form.
-Then the SOPs of $N$ SNN Conv layers, $M$ SNN Fully Connected Layer (FC), $L$ SSA , $F$ SSA and MWTF are added together are added together and multiplied by $E_{AC}$.
-The theoretical energy consumption for ANN and SNN models is calculated as follows:
-For ANNs, the energy consumption of a block $b$ is given by:
-$\text{Power}(b)=4.6\,\mathrm{pJ}\times\mathrm{FLOPs}(b).$ (35)
-For SNNs, the energy consumption of a block $b$ is given by:
-$\text{Power}(b)=0.9\,\mathrm{pJ}\times\mathrm{SOPs}(b).$ (36)
-V-F More detail of Ablation Studies
-In this section, we provide additional details regarding the ablation studies conducted in our research.
-These experiments are designed to investigate the contribution of various components within our framework and to demonstrate the effectiveness of specific design choices.
-By systematically removing or altering key elements, we aim to provide a deeper understanding of their impact on the overall performance.
-3D coordinate normalization
-3D skeleton data often spans multiple quadrants and includes negative values, as shown in Figure 2, creating challenges for SNNs, which are inherently designed for non-negative numerical inputs.
-Negative coordinates lead to significant information loss and degrade model performance.
-To address this limitation, 3D coordinate normalization is applied to map skeleton data into a fixed and bounded range, ensuring numerical consistency and preserving essential information.
-Normalization facilitates stable data propagation through the network by mitigating discrepancies arising from varying magnitudes and signs.
-Figure 2: 3D Skeleton Visualization
-The analysis focuses primarily on accuracy as the evaluation metric, as changes in parameters and FLOPs introduced by normalization are minimal.
-By isolating the impact of 3D coordinate normalization on accuracy, the study demonstrates that normalization significantly enhances model performance without introducing substantial computational overhead.
-Controlled experiments under various module configurations reveal consistent accuracy improvements, underscoring this approach’s effectiveness in processing complex skeleton data.
-Detailed results presented in the accompanying table highlight the practical benefits of normalization in achieving higher accuracy with negligible computational trade-offs.
-TABLE VI: Impact of 3D Coordinate Normalization on Accuracy
-3D Coord. Norm. 1D-SGC FSC MWTF Acc (%)
-- - - 69.7
-- - 77.0
-- 79.3
-- - 70.5
-- 78.3
-80.5
-Figure 3: Comparison of confusion matrices for full-module integration setup with and without 3D coordinate normalization
-The results in Table 3 demonstrate the significant role of 3D coordinate normalization in improving accuracy across all module configurations.
-For the baseline configuration (1D-SGC only), normalization increases accuracy by 0.8%, while in more complex setups with FSC and MWTF modules, the improvements reach 1.3% and 1.2%, respectively.
-This indicates that normalization effectively stabilizes the input data by addressing inconsistencies in coordinate ranges, enabling the model to better utilize its capacity.
-Importantly, the improvements in accuracy are achieved without notable changes in parameters or computational cost, highlighting normalization as a practical and efficient enhancement for processing 3D skeleton data.
-To further analyze the impact of 3D coordinate normalization, we selected the full-module integration setup (MWTF enabled) from the third and sixth rows of Table VI and visualized the corresponding confusion matrices.
-Different Settings for MWTF
-To investigate the impact of different configurations of $k$ (frequency resolution) and $J$ (decomposition levels) on the performance of the MWTF module within Signal-SGN, we conduct a series of controlled experiments.
-The overall architecture of Signal-SGN is kept unchanged, ensuring that any observed variations in performance can be attributed solely to changes in the MWTF parameters.
-By systematically adjusting $k$ and $J$, we aim to analyze how the trade-off between frequency granularity and decomposition depth affects the network’s ability to extract and fuse temporal features effectively.
-TABLE VII: Comparison of experimental results under different MWTF settings ($k$ and $J$)
-MWTF $k$ $J$ Param. (M) FLOPs/SOPs (G) Acc (%)
-- - - 1.65 1.60/0.312 78.3
-4 1 1.68 1.61/0.313 78.9
-8 1 1.73 1.62/0.314 79.8
-8 2 1.74 1.62/0.315 80.3
-8 3 1.74 1.62/0.317 80.5
-16 3 1.83 1.64/0.320 80.5
-The experimental results in Table VII highlight the impact of varying $k$ (frequency resolution) and $J$ (decomposition levels) on the performance of the MWTF module within Signal-SGN.
-With the overall architecture unchanged, the table isolates the influence of MWTF parameters, providing insight into their effects on accuracy, computational complexity, and parameter efficiency.
-When MWTF is disabled, the model achieves a baseline accuracy of 78.3% with a parameter count of 1.65M and computational costs of 1.60/0.312 G (FLOPs/SOPs).
-Enabling MWTF with different $k$ and $J$ configurations demonstrates significant improvements.
-As $k$ increases from $4$ to $8$, accuracy improves substantially, with the model achieving 79.8% ($k=8,J=1$).
-Further increasing $k$ to $16$ yields diminishing returns, with accuracy peaking at 80.5% for $k=16,J=3$.
-Similarly, increasing $J$ allows for deeper decomposition of temporal features, which enhances performance up to a point.
-For example, with $k=8$, accuracy improves from 79.8% ($J=1$) to 80.7% ($J=3$).
-This configuration achieves the best accuracy while maintaining a parameter count of 1.74M and computational cost of 1.62/0.317 G.
-Further increasing $k$ to $16$ ($J=3$) slightly reduces accuracy to 80.5%, indicating that excessively high frequency resolution or decomposition depth might lead to redundant features or overfitting.
-Overall, the results demonstrate that the MWTF module significantly enhances the performance of Signal-SGN when configured optimally.
-The configuration $k=8,J=3$ emerges as the most effective setting, balancing frequency granularity, decomposition depth, and computational efficiency, and achieving the best accuracy of 80.7%.
-This underscores the importance of selecting appropriate MWTF parameters to maximize feature extraction and model performance.
+The complete iterative decomposition process is outlined in Algorithm 2, providing a systematic and computationally efficient method for processing spike-form feature.
+Appendix D Supplementary Analysis
+D.1. Analysis of Window function
+(a) Window weights at initialization.
+(b) Window weights after training.
+Figure 8. Visualization of the learned joint-wise window functions $w_{V}(v)$ across FSC layers.
+To analyze the role of the learnable joint-wise window function $w_{V}(v)$, we visualize its values across all FSC layers at both the beginning and end of training, as shown in Figure 8.
+Initially, all layers exhibit near-zero or uniform weights, especially in deeper layers like Layer 4, where the window is almost flat and close to zero, as shown in Figure 8(a).
+This indicates that the network starts without any strong spatial preference in the frequency modulation stage.
+After training, however, the learned windows evolve into highly diverse and non-uniform patterns across joints,as shown in Figure 8(b).
+The early layers (e.g., Layer 1) show sharp variations, possibly focusing on localized joints, while deeper layers develop smoother yet still distinctive profiles.
+This progression suggests that learning a fully adaptive window function allows the network to emphasize or suppress specific joints based on spatiotemporal dynamics, with joint-wise spectral weighting acting as a spatial prior to guide frequency decomposition according to human motion structure and semantics.
+D.2. Analysis of Learnable Fusion Parameter $\beta$
+Figure 9. Joint evolution of training loss and the learnable fusion weight $\beta$.
+To complement the ablation study on the MWTF module, we provide an analysis of the learnable fusion weight $\beta$ in the final classification stage, as defined in Equation (17).
+Figure 9 illustrates the joint evolution of the training loss and the value of $\beta$ throughout training.
+We observe that $\beta$ fluctuates significantly during the initial 20 epochs, reflecting the model’s attempt to dynamically balance the temporal backbone output and the spectral feature $\hat{X}$ extracted by MWTF.
+As the training loss steadily declines, $\beta$ gradually stabilizes.
+By the end of training, $\beta$ converges to approximately 0.42, indicating that the model assigns moderate but consistent weight to the spectral branch.
+This behavior suggests that while the primary classification signal originates from the SN-processed temporal pathway, the MWTF-derived spectral features provide substantial complementary information.
+The learnable fusion mechanism, governed by $\beta$, thus plays a key role in adaptively integrating multiscale dynamics for robust action recognition.
+D.3. Frequency-Domain Analysis of MWTF
+To further interpret the effectiveness of the proposed spiking cross-attention fusion across wavelet decomposition levels, we perform a joint-wise spectral analysis using the example of the ”drinking water” action on NTU-RGB+D.
+As shown in Figure 10, each subfigure presents the magnitude spectrogram obtained from the FFT of a single channel with the highest temporal variance.
+The final cross-attentively fused output $\hat{X}$ in Figure 10(a) exhibits an enriched spectrum that combines both low- and high-frequency components.
+Compared to the low-frequency-dominated baseline $X^{(L)}$ in Figure 10(b), $\hat{X}$ demonstrates enhanced high-frequency responses particularly at joints 6 (right shoulder), 8 (right elbow), 10 (right wrist), and 12 (right hand), which are anatomically responsible for raising the arm and bringing the hand to the mouth during drinking.
+Figures 10(c)–(f) illustrate the spiking cross-attentively fused detail coefficients $\hat{D}_{0}$ to $\hat{D}_{3}$, representing progressively higher frequency bands as the decomposition level increases.
+Specifically, $\hat{D}_{0}$ and $\hat{D}_{1}$ capture mid-frequency activations in joints 6, 8, and 10, reflecting shoulder and elbow articulation during arm lifting.
+The component $\hat{D}_{2}$ highlights more localized and rapid changes in joints 10 and 12, aligning with the wrist-to-mouth trajectory.
+As the final decomposition level, $\hat{D}_{3}$ is generated via upsampling of the residual high-frequency component and retains only coarse responses—mainly from trunk-related joints (e.g., joints 2 and 3)—representing global pose adjustments.
+Figures 10(g)–(i) show the corresponding scaling coefficients $\hat{S}_{0}$ to $\hat{S}_{2}$, also fused through spiking cross-attention (cf.
+Eq.
+15).
+These components preserve low-frequency temporal dynamics and body stability.
+For example, $\hat{S}_{1}$ and $\hat{S}_{2}$ maintain dominant low-frequency energy at joints 2 (chest) and 3 (neck), which are critical for torso alignment and posture maintenance during the action.
+Overall, this frequency-domain decomposition provides fine-grained interpretability of motion dynamics across joint hierarchies and temporal scales, confirming that the proposed spiking cross-attention fusion mechanism effectively preserves both semantic structure and dynamic variations by adaptively emphasizing functionally critical joints in the spectral domain.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(a) $\hat{X}$ (b) $X^{(L)}$ (c) $\hat{D}_{0}$ (d) $\hat{D}_{1}$ (e) $\hat{D}_{2}$
+Refer to caption Refer to caption Refer to caption Refer to caption
+(f) $\hat{S}_{0}$ (g) $\hat{S}_{1}$ (h) $\hat{S}_{2}$ (i) $\hat{S}_{3}$
+Figure 10.
+Joint-wise frequency spectrograms of multiscale components, based on the ”drinking water” action.
+Each plot shows the FFT magnitude curves of all 25 joints, revealing their spectral activity at various decomposition levels.